### v1

# Homogeneous $\mathrm{G}_{2}$ and Sasakian instantons on the Stiefel $7$-manifold

Abstract
We study homogeneous instantons on the seven dimensional Stiefel manifold $V^{5,2}$ in the context of $\mathrm{G}_{2}$ and Sasakian geometry. According to the reductive decomposition of $V^{5,2}$, we provide an explicit description of all invariant $\mathrm{G}_{2}$ and Sasakian structures. In particular, we characterise the invariant $\mathrm{G}_{2}$-structures inducing a Sasakian metric, among which the well known nearly parallel $\mathrm{G}_{2}$-structure (Sasaki-Einstein) is included. As a consequence, we classify the invariant connections on homogeneous principal bundles over $V^{5,2}$ with gauge group ${\rm U}(1)$ and ${\rm SO}(3)$, satisfying either the $\mathrm{G}_{2}$ or the Sasakian instanton condition. Finally, we analyse the Yang Mills condition for those invariant connections.
1 Introduction
The Stiefel manifold $V^{n,k}$ of orthonormal $k$-frames in the $n$-dimensional Euclidean space (i.e. $1\leq k\leq n$) admits a transitive ${\rm SO}(n)$-action with isotropy subgroup isomorphic to ${\rm SO}(n-k)$, thus, it can be written homogeneously as $V^{n,k}={\rm SO}(n)/{\rm SO}(n-k)$ and consequently, its family of ${\rm SO}(n)$-invariant Riemannian metrics are described in terms of the isotropy representation of $V^{n,k}$. For instance, for $k=1$ or $k=n$, the isotropy representation is irreducible (indeed $V^{n,1}=S^{n}$ or $V^{n,n}={\rm SO}(n)$), in those cases, there is a unique ${\rm SO}(n)$-invariant metric, which is Einstein [wolf1968]. For $k=2$, the isotropy representation is reducible and the space of ${\rm SO}(n)$-invariant Riemannian metrics is described by a $4$-dimensional family, within there is a unique Einstein metric, which is inherited from the Grasmmanian $G_{2}(\mathbb{R}^{n})$ (e.g. [Arvanitoyeorgos1996, kerr1998new]). In addition, $V^{n,2}$ fits as an instance of the Boothby-Wang construction, as a total space of a $S^{1}$-principal bundle [boothby1958]*Theorem 3, moreover, it has a regular contact structure whose associated Reeb vector field generates a $1$-parameter subgroup of right translations of $V^{n,2}$.
A contact structure [^1] The reader unfamiliar with Sasakian geometry can find a complete study on this subject in [Boyer2008] on a $n$-dimensional manifold $M$ is a 1-form satisfying $\eta\wedge(d\eta)^{n}\neq 0$, in this case there exist a unique vector $\xi\in\mathscr{X}(M)$, called Reeb vector field, such that $\eta(\xi)=1$. The contact structure is called an almost contact structure if there exist an endomorphism $\Phi\in{\mathrm{End}}(TM)$ which satisfy $\Phi^{2}=-{\rm I}_{TM}+\eta\otimes\xi.$ A Riemannian metric $g$ is called compatible with the contact structure, if $g$ satisfies $g(\Phi X,\Phi Y)=g(X,Y)-\eta(X)\eta(Y),$ in this case, the contact structure is called an almost contact metric structure. In addition, if $\mathcal{S}=(\eta,\xi,g,\Phi)$ satisfies $2g(X,\Phi(Y))=d\eta(X,Y)$, the quadruple $\mathcal{S}$ is called a contact metric structure. Furthermore, if the Nijenhuis tensor $N_{\Phi}$ induced by $\Phi$ satisfies $N_{\Phi}=-d\eta\otimes\xi$, then $\mathcal{S}$ is called a Sasakian structure. In this paper, we consider the $7$–dimensional Stiefel manifold $V^{5,2}$ endowed with a one-parameter family of ${\rm SO}(5)$-invariant Sasakian structures $\mathcal{S}=(\eta,\xi,g,\Phi)$. For instance, homogeneous contact/Sasakian structures have been study in [andrada2009, correa2019] and [perrone1998homogeneous].
On the other hand, a differential $3$-form $\varphi$ on an oriented $7$–dimensional manifold $M$ is called a $\mathrm{G}_{2}$–structure if at each $p\in M$, there is a basis $e_{1},\dots,e_{7}$ of $T_{p}M$ such that
$\varphi_{p}=e^{125}+e^{136}+e^{147}+e^{234}-e^{267}+e^{357}-e^{456},$ (1.1)
where $e^{1},\dots,e^{7}$ denotes the dual basis of $T_{p}^{\ast}M$ and $e^{ijk}=e^{i}\wedge e^{j}\wedge e^{k}$. In particular, $\varphi$ determines a Riemannian metric $g_{\varphi}$, we study the family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$, such that the induced metric $g_{\varphi}$ is Sasakian. With respect to these structures, we study the theoretical Gauge instanton equations associated with the underlying Sasakian/$\mathrm{G}_{2}$-geometry of $V^{5,2}$. Namely, on a $G$–principal bundle $P\to M^{7}$ with $\sigma\in\Omega^{3}(M)$, the $1$-form connection $A\in\mathcal{A}(P)$ is called an $\lambda$-instanton, if it satisfies (e.g. [donaldson1998gauge, Tian2000])
$\ast(\sigma\wedge F_{A})=\lambda F_{A},\quad\text{where}\quad\lambda\in\mathbb{R}$ (1.2)
where $F_{A}=dA+A\wedge A$ denotes the corresponding curvature $2$-form. According to the specific pair $(\sigma,\lambda)$ in (1.2), we define:
  1. 1.
    A Self-dual contact Instantons (SDCI) [portilla2023] is a solutions of (1.2) with $\lambda=1$ and $\sigma:=\eta\wedge\omega$ naturally defined by a contact structure of $(M,\eta)$, where $\omega:=\frac{1}{2}d\eta$ is called the fundamental $2$-form..
  2. 2.
    A $\mathrm{G}_{2}$–instantons [corrigan1983first, donaldson1998gauge] is a solutions of (1.2) with $\lambda=-1$ and $\sigma=\varphi$ a coclosed (i.e. $d\ast\varphi=0$) $\mathrm{G}_{2}$–structure on $M$.
Furthermore, for $E\to(M^{7},\mathcal{S})$ a Sasakian holomorphic bundle [biswas2010vector], the usual notion of Hermitian Yang-Mills (HYM) connection on a Kähler manifolds is extended to odd dimension in the natural fashion to the so-called transverse Hermitian Yang-Mills connections (tHYM) (see (3.1)). The concept of Chern connections also extend naturally to a Sasakian version, as a connection mutually compatible with the transverse holomorphic structure and the Hermitian metric. A noteworthy result is that if $E\to M$ is a holomorphic Sasakian bundle on a contact Calabi-Yau manifold $M$, along compatible connections the notion of SDCI, $\mathrm{G}_{2}$-instantons and tHYM coincide. cf. [portilla2023, Teorem 1.1] and [calvo2020gauge, Lemma  21].
In [donaldson1998gauge, Donaldson2011], Donaldson, Segal and Thomas suggested that it may be possible to use $\mathrm{G}_{2}$-instantons to construct an enumerative invariant of $\mathrm{G}_{2}$-manifolds. For instance, when $(M^{7},\varphi)$ has a torsion-free $\mathrm{G}_{2}$-structure $\varphi$ (i.e. $d\varphi=0$ and $d\ast\varphi=0$ [fernandez1982riemannian]), solutions to the $\mathrm{G}_{2}$-instanton equation have been studied in [sa2015g2, walpuski2015g_2, clarke2014instantons, lotay20182]. However, the construction of metrics with holonomy $\mathrm{G}_{2}$ has proven to be a challenging task, thus, the torsion-free condition has been relaxed to the class of coclosed $\mathrm{G}_{2}$-structures [ball2019gauge, clarke2023], which has a notably application in theoretical physics to the Killing spinor equations in supergravity (cf. [Ivanov2005, Ivanov10, delaOssaG21, Lotay2022, Fernandez2011]). Analogously, Wang has proposed studying the compactification of the moduli space of contact instantons as a tool to construct a Donaldson-type invariant for contact $7$-manifolds [wang2012higher].
Description of the main results: Let $\mathfrak{so}(5)=\mathfrak{so}(3)\oplus{\mathfrak{m}}$ be a reductive decomposition of $V^{5,2}$, in [reidegeld2010spaces] Reidegeld proved that the space of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures has dimension $5$ and the space of induced ${\rm SO}(5)$-invariant Riemannian metrics has dimension $4$. Thus, we prove in Lemma 2.3 that the $\mathrm{Ad}({\rm SO}(3))$-invariant $3$-form (2.11) determined by $(a,b,x,y,z)\in\mathbb{R}^{5}$ defines a $\mathrm{G}_{2}$-structure (up to $\mathrm{G}_{2}$-transformation), if
$a^{2}+by>0,\quad b^{2}+ax>0\quad z\neq 0\quad\text{and}\quad xy=ab.$
In particular, it defines a family of (non-purely) coclosed $\mathrm{G}_{2}$-structures, within the nearly parallel $\mathrm{G}_{2}$-structures arises as an $S^{1}$-family for $x=a$, $y=b$ and $a^{2}+b^{2}=z^{2}=\frac{27}{512}$ (cf. Theorem 2.1).
In addition, for any $y_{1},y_{2},y_{3}\in\mathbb{R}^{+}$, Theorem 2.2 characterises the ${\rm SO}(5)$-invariant Sasakian metrics as
$y_{1}=4y_{2}^{2}\quad\text{and}\quad y_{2}=y_{3},$
as a consequence, for a given $(z,a,b,x,y)\in\mathbb{R}^{5}$, we describe the family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures inducing a Sasakian metric (cf. Corollary 2.8).
In Section 3, we revisit the interaction between SDCI, $\mathrm{G}_{2}$-instantons and tHYM connections addressed in [portilla2023, Proposition 3.11] and [calvo2020gauge, Lemma 21] against the backdrop of contact Calabi-Yau manifolds. For a bundle $E\to V^{5,2}$, in Lemma 3.1 we prove that if the connection $A$ is SDCI then it is a $\mathrm{G}_{2}$-instanton. Conversely, for a holomorphic Sasakian bundle $E\to V^{5,2}$, if the Chern connection is a $\mathrm{G}_{2}$-instanton, then it is a SCDI. Finally, we prove that HYM connections on a holomorphic bundle $E\to G_{2}(\mathbb{R}^{5})$ on the Grassmanian, pulling back to $\mathrm{G}_{2}$-instantons on $\pi^{\ast}E\to V^{5,2}$ (cf. Corollary 3.4). We recall here that the key argument follow just in presence of a SU$(3)$-structure.
Finally, in Section 4 we provide examples of homogeneous SCDI and $\mathrm{G}_{2}$-instantons over $V^{5,2}$, with gauge group ${\rm SO}(3)$ and ${\rm U}(1)$. In particular, we prove that the ${\rm SO}(3)$-invariant $\mathrm{G}_{2}$-instanton (4.8) is a Yang-Mills connection on $(V^{5,2},\varphi)$, for any ${\rm SO}(5)$-invariant coclosed $\mathrm{G}_{2}$-structure (cf. Theorem 4.1), moreover, it describes a local minimizer of the Yang-Mills functional (cf. Corollary 4.11).
Outline: In Section 2, we provide some preliminaries about $V^{5,2}$ as a homogeneous space, with the aim of characterising all the ${\rm SO}(5)$-invariant coclosed $\mathrm{G}_{2}$-structures and the associated Sasakian metrics. In Section 3, we explore the interaction between SDCI and $\mathrm{G}_{2}$–instantons on $V^{5,2}$, in particular, Lemmata 3.1, 3.3 and Corollary 3.4 indicate conditions such that contact instantons and $\mathrm{G}_{2}$-instantons coincides. Finally, in Section 4 we consider homogeneous principal bundles on $V^{5,2}$ and we construct examples of invariant instantons on ${\rm SO}(3)$, $U(1)$–homogeneous bundles. In particular, in Section 4.1 we characterise all invariant ${\rm SO}(3)$–connection and we show that these satisfy the $\mathrm{G}_{2}$ and SDCI condition for appropriate parameters and we analyse its stability as Yang-Mills connections.
Acknowledgments and funding: AM was funded by the São Paulo Research Foundation (Fapesp) [2021/08026-5] and [2023/13780-6]. LP was funded by Centre Henri Lebesgue (CHL) under the grant ANR-11-LABX-0020-01, he also would like to thanks to LMBA at Université de Bretagne Occidentale UBO for the hospitality.
2 Homogeneous $\mathrm{G}_{2}$ and contact metric structures on $V^{5,2}$
In this section, we provide an explicit description of a $\mathrm{Ad}({\rm SO}(3))$-invariant $\mathrm{G}_{2}$-structure and the induced inner product on $V^{5,2}$. As well as, we establish the notation that we will use throughout this work.
2.1 $\mathrm{G}_{2}$-structures on $V^{5,2}$
In order to describe the space of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structure on $V^{5,2}={\rm SO}(5)/{\rm SO}(3)$ [^2] This subject is brought up in [friedrich1997nearly]., we fix a reductive decomposition of $V^{5,2}$ as a homogeneous space. Consider the Lie algebra of ${\rm SO}(5)$
$\mathfrak{so}(5):=\{A\in\mathfrak{gl}(\mathbb{R}^{5})\;|\;A+A^{T}=0\},$ (2.1)
and denote by $E_{ij}$, the $5\times 5$ matrix with $1$ at the $(ij)$-entry and 00 elsewhere, and fix a basis for $\mathfrak{so}(5)$:
$\begin{split}e_{1}=&E_{12}-E_{21},\quad e_{4}=E_{15}-E_{51},\quad e_{7}=E_{25}-E_{52},\quad e_{10}=E_{54}-E_{45},\\ e_{2}=&E_{13}-E_{31},\quad e_{5}=E_{23}-E_{32},\quad e_{8}=E_{34}-E_{43},\quad\\ e_{3}=&E_{14}-E_{41},\quad e_{6}=E_{24}-E_{42},\quad e_{9}=E_{35}-E_{53}.\end{split}$ (2.2)
We consider the Lie algebra $\mathfrak{so}(3)$ embedded in $\mathfrak{so}(5)$ and it is spanned by $\{e_{8},e_{9},e_{10}\}$. The Lie bracket of $\mathfrak{so}(5)$ is given by the commutator describe in Table 1.
[,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] $e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $e_{6}$ $e_{7}$ $e_{8}$ $e_{9}$ $e_{10}$
$e_{1}$ 00 $-e_{5}$ $-e_{6}$ $-e_{7}$ $e_{2}$ $e_{3}$ $e_{4}$ 00 00 00
$e_{2}$ $e_{5}$ 00 $-e_{8}$ $-e_{9}$ $-e_{1}$ 00 00 $e_{3}$ $e_{4}$ 00
$e_{3}$ $e_{6}$ $e_{8}$ 00 $-e_{10}$ 00 $-e_{1}$ 00 $-e_{2}$ 00 $e_{4}$
$e_{4}$ $e_{7}$ $e_{9}$ $e_{10}$ 00 00 00 $-e_{1}$ 00 $-e_{2}$ $-e_{3}$
$e_{5}$ $-e_{2}$ $e_{1}$ 00 00 00 $-e_{8}$ $-e_{9}$ $e_{6}$ $e_{7}$ 00
$e_{6}$ $-e_{3}$ 00 $e_{1}$ 00 $e_{8}$ 00 $-e_{10}$ $-e_{5}$ 00 $e_{7}$
$e_{7}$ $-e_{4}$ 00 00 $e_{1}$ $e_{9}$ $e_{10}$ 00 00 $-e_{5}$ $-e_{6}$
$e_{8}$ 00 $-e_{3}$ $e_{2}$ 00 $-e_{6}$ $e_{5}$ 00 00 $-e_{10}$ $e_{9}$
$e_{9}$ 00 $-e_{4}$ 00 $e_{2}$ $-e_{7}$ 00 $e_{5}$ $e_{10}$ 00 $-e_{8}$
$e_{10}$ 00 00 $-e_{4}$ $e_{3}$ 00 $-e_{7}$ $e_{6}$ $-e_{9}$ $e_{8}$ 00
Table 1: Bracket of $\mathfrak{so}(5)$ with respect to the basis (2.2)
On $\mathfrak{so}(5)$, we fix the inner product $\langle a,b\rangle=Q(a,b)=\frac{1}{2}\mathop{\mathrm{Tr}}\nolimits(ab^{T}),$ with respect to which, we have the orthogonal splitting
$\mathfrak{so}(5)={\mathfrak{m}}\oplus\mathfrak{so}(3)\quad\text{where}\quad{\mathfrak{m}}:=\mathfrak{so}(3)^{\perp}=\operatorname{{\rm Span}}\{e_{1},...,e_{7}\}.$ (2.3)
Note that in terms of the basis (2.2), the restriction of $\mathrm{ad}(e_{k})_{{\mathfrak{m}}}$ to ${\mathfrak{m}}$ are given by
$\mathrm{ad}(e_{8})_{\mathfrak{m}}$ $=\mathrm{diag}(0,A_{1},A_{1}),$
$\mathrm{ad}(e_{9})_{\mathfrak{m}}$ $=\mathrm{diag}(0,A_{2},A_{2}),$
$\mathrm{ad}(e_{10})_{\mathfrak{m}}$ $=\mathrm{diag}(0,A_{3}A_{3}),$
where $A_{1}=\begin{pmatrix}0&-1&0\\ 1&0&0\\ 0&0&0\end{pmatrix}$,  $A_{2}=\begin{pmatrix}0&0&-1\\ 0&0&0\\ 1&0&0\end{pmatrix}$ and $A_{3}=\begin{pmatrix}0&0&0\\ 0&0&-1\\ 0&1&0\end{pmatrix}$ are the canonical basis of $\mathfrak{so}(3)$. Denote by ${\mathfrak{m}}_{1}=\operatorname{{\rm Span}}\{e_{1}\},$ ${\mathfrak{m}}_{2}:=\operatorname{{\rm Span}}\{e_{2},e_{3},e_{4}\}$ and ${\mathfrak{m}}_{3}:=\operatorname{{\rm Span}}\{e_{5},e_{6},e_{7}\}.$ By using the Lie bracket in Table 1 is straightforward to verify that
$\mathrm{ad}(\mathfrak{so}(3)){\mathfrak{m}}_{1}=0,\quad\mathrm{ad}(\mathfrak{so}(3)){\mathfrak{m}}_{2}\subseteq{\mathfrak{m}}_{2}\quad\text{and}\quad\mathrm{ad}(\mathfrak{so}(3)){\mathfrak{m}}_{3}\subseteq{\mathfrak{m}}_{3}.$
Consequently, for the reductive decomposition (2.3) we obtain the irreducible $\mathrm{ad}(\mathfrak{so}(3))$-submodules
${\mathfrak{m}}={\mathfrak{m}}_{1}\oplus{\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}.$ (2.4)
The subalgebra $\mathfrak{so}(2)\oplus\mathfrak{so}(3)=\operatorname{{\rm Span}}\{e_{1},e_{8},e_{9},e_{10}\}\subset\mathfrak{so}(5)$ provides a reductive decomposition $\mathfrak{so}(5)=\mathfrak{so}(2)\oplus\mathfrak{so}(3)\oplus{\mathfrak{p}}$ (where ${\mathfrak{p}}={\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}$) for the symmetric space
$G_{2}(\mathbb{R}^{5})=\frac{{\rm SO}(5)}{{\rm SO}(2)\times{\rm SO}(3)},$ (2.5)
furthermore, ${\mathfrak{p}}\simeq\mathbb{C}^{3}$ with respect to the canonical almost complex structure $J=\begin{pmatrix}0&-I_{3}\\ I_{3}&0\end{pmatrix}$ and it turns out that ${\mathfrak{m}}_{3}=\operatorname{{\rm Span}}\{Je_{2},Je_{3},Je_{4}\}$. With respect to the basis (2.2), we define the alternating forms on ${\mathfrak{p}}$ given by
$\begin{array}[]{rl}\omega_{0}&=e^{25}+e^{36}+e^{47},\\[3.0pt] \mathop{\mathrm{Im}}\Omega_{0}&=e^{237}-e^{246}+e^{345}-e^{567},\\[3.0pt] \mathop{\mathrm{Re}}\Omega_{0}&=e^{234}-e^{267}+e^{357}-e^{456}.\end{array}$ (2.6)
Notice that $\omega_{0}\wedge\mathop{\mathrm{Re}}\Omega_{0}=0$ and $2\omega_{0}^{3}=3\mathop{\mathrm{Re}}\Omega_{0}\wedge\mathop{\mathrm{Im}}\Omega_{0}\neq 0$, i.e. $(\omega_{0},\mathop{\mathrm{Re}}\Omega_{0})$ is a ${\rm SU}(3)$-structure on ${\mathfrak{p}}$ (with induced orientation $-e^{234567}$).
Lemma 2.1.
The forms in (2.6) are $\mathrm{ad}(\mathfrak{so}(2)\oplus\mathfrak{so}(3))$-invariant, therefore, it induces a ${\rm SO}(5)$-invariant ${\rm SU}(3)$-structure $(\omega,\mathop{\mathrm{Re}}\Omega)$ on $G_{2}(\mathbb{R}^{5})$. Furthermore, we have that $d\omega=0,$ $d(\mathop{\mathrm{Re}}\Omega)=3e^{1}\wedge\mathop{\mathrm{Im}}\Omega$ and $d(\mathop{\mathrm{Im}}\Omega)=-3e^{1}\wedge\mathop{\mathrm{Re}}\Omega.$
Proof.
The Lie algebras $\mathfrak{sl}(\mathbb{C}^{3})$ and ${\mathfrak{s}}{\mathfrak{p}}(\mathbb{R}^{6})$ are described by:
$\mathfrak{sl}(\mathbb{C}^{3})=\left\{\begin{pmatrix}B&-C\\ C&B\end{pmatrix}\in\mathfrak{gl}({\mathfrak{p}})\Big{|}B,C\in\mathfrak{sl}(\mathbb{R}^{3})\right\},\quad{\mathfrak{s}}{\mathfrak{p}}(\mathbb{R}^{6})=\left\{\begin{pmatrix}B&C\\ D&-B^{T}\end{pmatrix}\in\mathfrak{gl}({\mathfrak{p}})\Big{|}C,D\in sym(\mathbb{R}^{3})\right\}.$
The Lie algebras $\mathfrak{sl}(\mathbb{C}^{3})$ and ${\mathfrak{s}}{\mathfrak{p}}(\mathbb{R}^{6})$ can be interpreted as stabilisers
$\mathfrak{stab}_{\mathfrak{gl}({\mathfrak{p}})}(\omega_{0})=\mathfrak{sl}(\mathbb{C}^{3})$
and
$\mathfrak{stab}_{\mathfrak{gl}({\mathfrak{p}})}(\mathop{\mathrm{Re}}\Omega_{0})=\mathfrak{stab}_{\mathfrak{gl}({\mathfrak{p}})}(\mathop{\mathrm{Im}}\Omega_{0})=\mathfrak{sp}(\mathbb{R}^{6}).$
Notice that the endomorphisms $\mathrm{ad}(e_{8})_{\mathfrak{p}}$, $\mathrm{ad}(e_{9})_{\mathfrak{p}}$, $\mathrm{ad}(e_{10})_{\mathfrak{p}}$ and $\mathrm{ad}(e_{1})_{\mathfrak{p}}=J$ belong to the interception $\mathfrak{sl}(\mathbb{C}^{3})\cap\mathfrak{sp}(\mathbb{R}^{6})$. Therefore, the forms in (2.6) are $\mathrm{ad}(\mathfrak{so}(2)\oplus\mathfrak{so}(3))$-invariant. Now, for any $p=x\cdot\Big{(}{\rm SO}(2)\times{\rm SO}(3)\Big{)}\in G_{5,2}$ define $(\omega,\real\Omega)$ by
$\omega_{p}={\rm L}_{x}^{\ast}\omega_{0}\quad\text{and}\quad\real\Omega_{p}={\rm L}_{x}^{\ast}\real\Omega_{0},$
where ${\rm L}_{x}$ is the left multiplication by $x\in{\rm SO}(5)$, hence $2\omega^{3}=3\mathop{\mathrm{Re}}\Omega\wedge\mathop{\mathrm{Im}}\Omega\neq 0$ Finally, from Table 1 follows that $d\omega_{0}=0,$ $d(\mathop{\mathrm{Re}}\Omega_{0})=3e^{1}\wedge\mathop{\mathrm{Im}}\Omega_{0}$ and $d(\mathop{\mathrm{Im}}\Omega_{0})=-3e^{1}\wedge\mathop{\mathrm{Re}}\Omega_{0}$. ∎
The following result is mentioned in [kerr1998new]*§4, for sake of completeness, we include the detail proof of it.
Lemma 2.2.
Any $\mathrm{Ad}({\rm SO}(3))$-invariant metric on $V^{5,2}$ is isometric to a diagonal one, i.e., up to isometries, it is enough to consider the class of invariant metrics determined by the $\mathrm{Ad}({\rm SO}(3))$-invariant inner product:
$g=y_{1}Q|_{{\mathfrak{m}}_{1}}+y_{2}Q|_{{\mathfrak{m}}_{2}}+y_{3}Q|_{{\mathfrak{m}}_{3}},\quad\text{where}\quad y_{i}>0\quad i=1,2,3.$ (2.7)
Proof.
In the decomposition (2.4) we have ${\mathfrak{m}}_{2}\simeq{\mathfrak{m}}_{3}$, hence, any ${\rm SO}(5)$-invariant metric on $V^{5,2}$ is determined by the $\mathrm{Ad}({\rm SO}(3))$-invariant inner product $\langle\cdot,\cdot\rangle=Q(g\cdot,\cdot)$, where $g$ is a positive symmetric matrix of the form
$g=\begin{pmatrix}y_{1}&0&0\\ 0&y_{2}I_{3}&xI_{3}\\ 0&xI_{3}&y_{3}I_{3}\end{pmatrix}\quad\text{for}\quad y_{1},y_{2},y_{3}>0,\quad x\in\mathbb{R}\quad\text{with}\quad y_{2}y_{3}>x^{2}.$
Denote by $N({\rm SO}(3))$ the normalizer of ${\rm SO}(3)$ in ${\rm SO}(5)$, since the quotient $\frac{N({\rm SO}(3))}{{\rm SO}(3)}$ is isomorphic to ${\rm SO}(2)$ the conjugation induces a 1-parameter family of diffeomorphism on ${\mathfrak{m}}$ given by $\phi(t)=\mathrm{Ad}(\exp(te_{1}))$. Thus, for $t\in\mathbb{R}$ the one-parameter family $g_{t}=\phi(t)^{T}g\phi(t)$ defines an isometric $\mathrm{Ad}({\rm SO}(3))$-inner product.
A straightforward computation shows that the parameter $x(t)$ in the metric $g_{t}$, is given by $x(t)=\frac{(y_{3}-y_{2})}{2}\sin(2t)+x\cos(2t)$. Hence, $g_{t}$ is diagonal if:
$y_{2}\neq y_{3}\quad\text{and}\quad t=\frac{1}{2}\arctan\left(\frac{2x}{y_{2}-y_{3}}\right)$
or
$y_{2}=y_{3}\quad\text{and}\quad t=\frac{(2k+1)\pi}{4}\quad\text{with}\quad k\in{\bf N},$
With respect to (2.7), we define an orthonormal basis of ${\mathfrak{m}}$:
$X_{1}:=\frac{1}{\sqrt{y_{1}}}e_{1},\quad X_{i}:=\frac{1}{\sqrt{y_{2}}}e_{i}\quad\text{and}\quad X_{i+3}:=\frac{1}{\sqrt{y_{3}}}e_{i+3}\quad\text{for}\quad i=1,2,3.$ (2.8)
Then, the Lie bracket in Table 1 becomes
$\begin{split}[X_{1},X_{i}]_{{\mathfrak{m}}}=-\sqrt{\frac{y_{3}}{y_{1}y_{2}}}X_{i+3},\quad[X_{1},X_{i+3}]_{{\mathfrak{m}}}=\sqrt{\frac{y_{2}}{y_{1}y_{3}}}X_{i},\quad[X_{i},X_{i+3}]_{{\mathfrak{m}}}=-\sqrt{\frac{y_{1}}{y_{2}y_{3}}}X_{1}.\end{split}$ (2.9)
In [reidegeld2010spaces], Reidegeld proved that the space of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$ has dimension $5$. The next Lemma describes the invariant $\mathrm{G}_{2}$-structures on the Stiefel manifold up to $\mathrm{G}_{2}$-equivalence, i.e. for a $\mathrm{G}_{2}$-structure $\varphi$ written in the basis $\{e_{1},...,e_{7}\}$ and $\{\tilde{e}_{1},...,\tilde{e}_{7}\}$ of ${\mathfrak{p}}$, there is a change of basis $f\in\mathrm{G}_{2}$ from $\{e_{1},...,e_{7}\}$ to $\{\tilde{e}_{1},...,\tilde{e}_{7}\}$.
Lemma 2.3.
The $5$-dimensional family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$ are described, up to $\mathrm{G}_{2}$-equivalence, by $(a,b,x,y,z)\in\mathbb{R}^{5}$ satisfying
$a^{2}+by>0,\quad b^{2}+ax>0\quad z\neq 0\quad\text{and}\quad xy=ab.$ (2.10)
Explicitly, the $\mathrm{G}_{2}$-structure $\varphi$ can be written by
$\varphi=$ $-X^{1}\wedge(X^{25}+X^{36}+X^{47})+p_{1}X^{234}+p_{2}(-X^{267}+X^{357}-X^{456})$ (2.11)
$+p_{3}(X^{237}-X^{246}+X^{345})-p_{4}X^{567},$
where
$p_{1}=\frac{x\sqrt{a^{2}+by}}{b^{2}+ax},\quad p_{2}=\frac{a}{\sqrt{a^{2}+by}},\quad p_{3}=\frac{b}{\sqrt{b^{2}+ax}}\quad\text{and}\quad p_{4}=\frac{y\sqrt{b^{2}+ax}}{a^{2}+by}$ (2.12)
and
$y_{1}=\frac{z^{2}}{y_{2}y_{3}},\quad y_{2}=\frac{(b^{2}+ax)^{2/3}}{(a^{2}+by)^{1/3}}\quad\text{and}\quad y_{3}=\frac{(a^{2}+by)^{2/3}}{(b^{2}+ax)^{1/3}}.$ (2.13)
Proof.
From (2.6) is easy to check that any $\mathrm{Ad}({\rm SO}(3))$-invariant $3$-form on ${\mathfrak{m}}$ is given by
$\widetilde{\varphi}=$ $-\widetilde{z}e^{1}\wedge(e^{25}+e^{36}+e^{47})+\widetilde{x}e^{234}+\widetilde{a}(-e^{267}+e^{357}-e^{456})$
$+\widetilde{b}(e^{237}-e^{246}+e^{345})-\widetilde{y}e^{567}.$
The $3$-form $\widetilde{\varphi}$ defines a $\mathrm{G}_{2}$-structure, if the induced symmetric bilinear form
$B_{ij}=\Big{(}(e_{i}\lrcorner\varphi)\wedge(e_{j}\lrcorner\varphi)\wedge\varphi\Big{)}(e_{1},...,e_{7})$
is positive (or negative) defined. Explicitly, we obtain that for $k=2,3,4$
$B_{11}=6\widetilde{z}^{3},\quad B_{kk}=6\widetilde{z}(\widetilde{b}^{2}+\widetilde{a}\widetilde{x}),\quad B_{k+3k+3}=6\widetilde{z}(\widetilde{a}^{2}+\widetilde{b}\widetilde{x}),\quad B_{kk+3}=3\widetilde{z}(\widetilde{a}\widetilde{b}-\widetilde{x}\widetilde{y})$
and $B_{ij}=0$ otherwise.
Now, since $\phi(t)=\mathrm{Ad}(\exp(te_{1}))\in{\rm SU}(3)\subset\mathrm{G}_{2}$ [cf. Lemma 2.2], there exist $t_{0}\in\mathbb{R}$ such that
$\varphi=\phi(t_{0})^{*}\widetilde{\varphi}=$ $-ze^{1}\wedge(e^{25}+e^{36}+e^{47})+xe^{234}+a(-e^{267}+e^{357}-e^{456})$
$+b(e^{237}-e^{246}+e^{345})-ye^{567},$
satisfies (2.10). On the other hand, the induced metric is given by
$g=\frac{B}{6^{\frac{2}{9}}\mathop{\mathrm{missing}}{det}\nolimits(B)^{\frac{1}{9}}},$
hence, by comparing $g$ with (2.7), we obtain the condition in (2.13). Finally, the expression in (2.11) is obtained by writing $\varphi$ in the orthonormal basis (2.8). ∎
Definition 2.4.
A $\mathrm{G}_{2}$-structure $\varphi$ is called coclosed if $d\psi=0$. Among the class of coclosed class, $\varphi$ is said purely coclosed if $d\varphi\wedge\varphi=0$ and nearly parallel if $d\varphi=4\psi$.
From Lemma (2.3), the following result describe all coclosed invariant $\mathrm{G}_{2}$–metrics on $V^{5,2}$
Theorem 2.1.
The $\mathrm{G}_{2}$-structure (2.11) is coclosed for any $(a,b,x,y,z)\in\mathbb{R}^{5}$ satisfying (2.10). Moreover, $V^{5,2}$ does not carry a ${\rm SO}(5)$-invariant purely coclosed $\mathrm{G}_{2}$-structure and (2.11) is nearly parallel if
$x=a,\quad y=b,\quad a^{2}+b^{2}=\frac{27}{512}\quad\text{and}\quad z=-\frac{9}{32}.$ (2.14)
Proof.
From Lemma 2.3, we get the dual 4-form $\psi=\ast\varphi$
$\psi=$ $-(X^{2356}+X^{2457}+X^{3467})-p_{4}X^{1234}+p_{3}X^{1}\wedge(X^{267}-X^{357}+X^{456})$ (2.15)
$+p_{2}X^{1}\wedge(X^{237}-X^{246}+X^{345})-p_{1}X^{1567}.$
Using the Lie bracket (2.9) and the metric coefficients (2.7), we obtain the exterior derivative for the following $\mathrm{Ad}({\rm SO}(3))$-forms on ${\mathfrak{m}}$:
$dX^{1}=$ $\frac{z}{y_{2}y_{3}}\left(X^{25}+X^{36}+X^{47}\right)$ (2.16)
$dX^{234}=$ $-\frac{y_{2}}{z}\left(X^{1237}-X^{1246}+X^{1345}\right),\quad dX^{567}=\frac{y_{3}}{z}\left(X^{1267}-X^{1357}+X^{1456}\right)$
$d(X^{267}$ $-X^{357}+X^{456})=\frac{2y_{3}}{z}\left(X^{1237}-X^{1246}+X^{1345}\right)-\frac{3y_{2}}{z}X^{1567}$
$d(X^{237}$ $-X^{246}+X^{345})=\frac{3y_{3}}{z}X^{1234}-\frac{2y_{2}}{z}\left(X^{1267}-X^{1357}+X^{1456}\right).$
From (2.15) and (2.16), it is easy to check that $d\psi=0$. Now, from (2.11) we obtain
$d\varphi=$ $\frac{2z}{y_{2}y_{3}}\left(X^{2356}+X^{2457}+X^{3467}\right)-\frac{p_{1}y_{2}+2p_{2}y_{3}}{z}\left(X^{1237}-X^{1246}+X^{1345}\right)$ (2.17)
$+\frac{3p_{2}y_{2}}{z}X^{1567}+\frac{3p_{3}y_{3}}{z}X^{1234}-\frac{p_{4}y_{3}+2p_{3}y_{2}}{z}\left(X^{1267}-X^{1357}+X^{1456}\right).$
From a straightforward computation, we note that $\ast\left(d\varphi\wedge\varphi\right)=12+\frac{3}{y_{2}}+\frac{3}{y_{3}}\neq 0,$ in other words, $\varphi$ given in (2.11) is not purely coclosed. Finally, by imposing the condition $d\varphi=4\psi$, and using (2.15) and (2.17), we obtain (2.14). ∎
Remark 2.5.
The $S^{1}$-family given in (2.14) correspond to the general construction given [alexandrov2012]*(4.28).
2.2 Homogeneous Sasakian metrics on $V^{5,2}$
In this section, we study the Sasakian condition for ${\rm SO}(5)$-invariant metrics on $V^{5,2}$. Firstly, we characterise the class $(\eta,\xi,g,\Phi)$ of contact metrics structures, cf. Section 1. Secondly, in Theorem 2.2, we provide the constraints in the metric to get a Sasakian structures. And finally, we relate the invariant $\mathrm{G}_{2}$-structures inducing a Sasakian metric.
Lemma 2.6.
The globally defined $1$-form $\eta=X^{1}$ and $\xi=X_{1}$ define a ${\rm SO}(5)$-invariant almost contact structure on $V^{5,2}$. Furthermore, $(\eta,\xi,g,\Phi)$ defines a contact metric structure on $V^{5,2}$, if and only if
$g=4y_{2}y_{3}Q|_{{\mathfrak{m}}_{1}}+y_{2}Q|_{{\mathfrak{m}}_{2}}+y_{3}Q|_{{\mathfrak{m}}_{3}}\quad\text{and}\quad\Phi(X)=[e_{1},X].$ (2.18)
Proof.
First, note that $(V^{5,2},\eta)$ is a contact manifold, since it is clear that $\eta(\xi)=1$ and that $\eta\wedge d\eta^{3}\neq 0$ follows from (2.16). Now, we define an $\mathrm{Ad}({\rm SO}(3))$-invariant contact metric structure $(\eta,\xi,g,\Phi)$, i.e., we define an $\mathrm{Ad}({\rm SO}(3)$-invariant endomorphism $\Phi$ on ${\mathfrak{m}}$ such that $\Phi^{2}=-I+\eta\otimes\xi$ and $d\eta(X,Y)=2g(X,\Phi(Y))$. From (2.9) we have
$d\eta=\sqrt{\frac{y_{1}}{y_{2}y_{3}}}\left(X^{25}+X^{36}+X^{47}\right),$
Hence, the condition $d\eta(X_{i},X_{i+3})=2g(X_{i},\Phi(X_{i+3}))$ implies that $\Phi$ is given by
$\Phi X_{i}=-\frac{1}{2}\sqrt{\frac{y_{1}}{y_{2}y_{3}}}X_{i+3},\quad\Phi X_{i+3}=\frac{1}{2}\sqrt{\frac{y_{1}}{y_{2}y_{3}}}X_{i},\quad\text{and}\quad\Phi\xi=0.$ (2.19)
Moreover, using the above equalities and the condition $\Phi^{2}=-I+\eta\otimes\xi$, we obtain that $-\frac{1}{4}\frac{y_{1}}{y_{2}y_{3}}=-1$, or equivalently $y_{1}=4y_{2}y_{3}$. ∎
The Nijenhuis tensor induced by $\Phi$ is defined by
$N_{\Phi}(X,Y)=[\Phi X,\Phi Y]+\Phi^{2}[X,Y]-\Phi[\Phi X,Y]-\Phi[X,\Phi Y].$ (2.20)
From [Boyer2008]*Theorem 6.5.9, we know that an almost contact structure $(\xi,\eta,\Phi)$ is normal if the Nijenhuis tensor satisfies
$N_{\Phi}=-d\eta\otimes\xi.$ (2.21)
Following [Boyer2008]*Definition 6.5.13, we define:
Definition 2.7.
A normal contact metric structure $\mathcal{S}=(\eta,\xi,g,\Phi)$ on $M^{2n+1}$ is called Sasakian and $(M^{2n+1},\mathcal{S})$ is called a Sasakian manifold.
Theorem 2.2.
The ${\rm SO}(5)$-invariant contact metric structure $(\eta,\xi,g,\Phi)$ given in Lemma 2.6 is a Sasakian structure, if and only if, the metric (2.18) satisfies
$y_{1}=4y_{2}^{2}\quad\text{and}\quad y_{2}=y_{3}.$ (2.22)
Proof.
We show that $(\eta,\xi,g,\Phi)$ is a normal, contact metric structure, i.e., Sasakian. Using the bracket (2.9) and the definition of $\Phi$ in (2.18), we verify that the Nijenhuis tensor (2.20) satisfies (2.21). From Lemma 2.6, we recall that the metric $g$ is contact for $y_{1}=4y_{2}y_{3}$. First, note that $N_{\Phi}(\xi,Y)=\Phi^{2}[\xi,Y]-\Phi[\xi,\Phi Y]$ and now, for $i=1,2,3$ we consider
$\Phi^{2}[\xi,X_{i}]-\Phi[\xi,\Phi X_{i}]$ $=-[\xi,X_{i}]-\Phi[\xi,-X_{i+3}]=-[\xi,X_{i}]+\sqrt{\frac{y_{2}}{y_{1}y_{3}}}\Phi X_{i}$
$=\sqrt{\frac{y_{3}}{y_{1}y_{2}}}X_{i+3}-\sqrt{\frac{y_{2}}{y_{1}y_{3}}}X_{i+3}$
$=\left(\frac{1}{2y_{2}}-\frac{1}{2y_{3}}\right)X_{i+3}$
Since $i_{\xi}d\eta=0$, we conclude that $y_{2}=y_{3}$. Analogously
$\Phi^{2}[\xi,X_{i+3}]-\Phi[\xi,\Phi X_{i+3}]$ $=-[\xi,X_{i+3}]-\Phi[\xi,X_{i}]=\sqrt{\frac{y_{2}}{y_{1}y_{3}}}X_{i}-\sqrt{\frac{y_{3}}{y_{1}y_{2}}}\Phi X_{i+3}$
$=\frac{1}{\sqrt{y_{1}}}X_{i}-\frac{1}{\sqrt{y_{1}}}X_{i}=0.$
For the missing cases, for $X,Y\in{\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}$ the brackets $[X,Y]$, $[\Phi Y,Y]$ and $[X,\Phi Y]$ are generated by $\xi$ (cf. (2.9)), hence $\Phi[X,Y]=\Phi[\Phi X,Y]=\Phi[X,\Phi Y]=0$ and thus $N_{\Phi}(X,Y)=[\Phi X,\Phi Y]$. So, (2.21) holds by noting that
$N_{\Phi}(X_{i},X_{i+3})=-2X_{1}\quad\text{for}\quad i=2,3,4\quad\text{and}\quad N_{\Phi}(X_{i},X_{j})=0\quad\text{otherwise}.$
According to the notation in Theorem 2.2, the Sasakian metric (2.7) is induced by the following family of $\mathrm{G}_{2}$-structures.
Corollary 2.8.
Consider the Sasakian metric on $V^{5,2}$ as a metric induced by (2.11) and the parameter vector $(z,a,b,x,y)\in\mathbb{R}^{5}$ satisfying the condition (2.10). Then, the $\mathrm{G}_{2}$-structure (2.11) induces a Sasakian metric, if and only if,
$x=a,\quad y=b\quad\text{and}\quad z^{2}=4(a^{2}+b^{2})^{\frac{4}{3}}.$ (2.23)
In this case, the $2$-parameter family of coclosed $\mathrm{G}_{2}$-structures is
$\varphi=$ $-X^{1}\wedge(X^{25}+X^{36}+X^{47})+\left(\frac{2}{z}\right)^{\frac{3}{4}}\Big{\{}a\Big{(}X^{234}-X^{267}+X^{357}-X^{456}\Big{)}$
$+b\Big{(}X^{237}-X^{246}+X^{345}-X^{567}\Big{)}\Big{\}},$
$\psi=$ $-\Big{(}X^{2356}+X^{2457}+X^{3467}\Big{)}-\left(\frac{2}{z}\right)^{\frac{3}{4}}\Big{\{}bX^{1}\wedge\Big{(}X^{234}-X^{267}+X^{357}-X^{456}\Big{)}$
$-aX^{1}\wedge\Big{(}X^{237}-X^{246}+X^{345}-X^{567}\Big{)}\Big{\}}.$
Proof.
To proof item $(1)$, we compare (2.13) and (2.22), thus
$z^{2}=4(a^{2}+by)^{4/3}\quad\text{and}\quad a^{2}+by=b^{2}+ax>0,$
since $a^{2}b^{2}=abxy$, we have $a^{2}(ax+b^{2})=(b^{2}+ax)ax$ and $by(a^{2}+by)=b^{2}(by+a^{2})$. Thus, the result follows.
Remark 2.9.
On the $\mathrm{ad}(\mathfrak{so}(3))$-invariant reductive decomposition ${\mathfrak{m}}={\mathfrak{m}}_{1}\oplus{\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}$ (2.3) the tensors given by
$\xi=\frac{4}{3}e_{1},\quad\Phi=\begin{pmatrix}0&&\\ &&I_{3}\\ &-I_{3}&\end{pmatrix}\quad\text{and}\quad g=\frac{3}{8}\begin{pmatrix}\frac{3}{8}&0\\ 0&I_{6}\end{pmatrix},$
define the ${\rm SO}(5)$-invariant Sasakian-Einstein structure on $V^{5,2}$, associated with nearly parallel $\mathrm{G}_{2}$-structure of type II (cf.[friedrich1997nearly]*Table 2).
3 $\mathrm{G}_{2}$ and self dual contact instantons on $V^{5,2}$
According to Corollary 2.8, we can consider solutions of (1.2) for those $\mathrm{G}_{2}$-structures with induced Sasakian metric. Additionally, if we endow a complex vector bundle $E\to V^{5,2}$ with a holomorphic Sasakian structure [biswas2010vector], one obtain the notion of Chern connection, as those connections simultaneously compatible with the holomorphic structure and some Hermitian bundle metric. The Hermitian Yang-Mills (HYM) condition is naturally extended to a Sasakian version, namely, let $\omega\colon=d\eta\in\Omega^{1,1}(M)$ be the transversal Kähler form, we say that a connection $A\in\mathcal{A}(E)$ is transverse Hermitian Yang-Mills connection (tHYM) if
$\hat{F}_{A}\colon=\langle F_{A},\omega\rangle=0,\quad\text{and}\quad F^{0,2}_{A}=0.$ (3.1)
Here, the type $(p,q)$ of a $k$-form on $M$ is induced by the condition $\Phi|_{H}^{2}=-1$, where $H$ is the Kähler distribution $H=\ker\eta\subset TM=H\oplus\xi$. Additionally, the splitting on the tangent bundle induced by the contact structure yields the splitting $\Omega^{k}(M)=\Omega^{k}(H^{\ast})\oplus\eta\wedge\Omega^{k-1}(H^{\ast})$. Hence, after complexification of the tangent bundle, one can naturally extend this bi-degree $(p,q)$ to the bundle of $E$-valued $k$-forms. We know that $2$-forms on $M$, and consequently $\Omega^{2}({\mathfrak{g}}_{E})$, decomposes by (cf. [portilla2023, Section. 2.1])
$\Omega^{2}(M)=\underbrace{\Omega^{2}_{1}\oplus\Omega^{2}_{6}\oplus\Omega^{2}_{8}}_{\text{Transversal $2$-forms}}\oplus\Omega^{2}_{V}$ (3.2)
Furthermore, a connection $A\in\mathcal{A}(E)$ is SDCI if its curvature belong to $\Omega^{2}_{8}({\mathfrak{g}}_{E})$, or equivalently, if the curvature $F_{A}\in\Omega^{2}({\mathfrak{g}}_{E})$ is of type $(1,1)$ orthogonal to $\omega$ (cf. [portilla2023, Proposition 2.5]).
Lemma 3.1.
With respect to the Sasakian structure of $V^{5,2}$ given in Theorem 2.2 and the $\mathrm{G}_{2}$-structure in (2.11) we have that: If $A$ is a SDCI, then it is a $\mathrm{G}_{2}$-instanton.
Proof.
We know that in local coordinates, a basis for self dual contact $2$–forms is given by (cf. [portilla2023, Section 2])
$\begin{array}[]{lccl}w_{1}=X^{23}+X^{56},&w_{2}=X^{26}+X^{35},&w_{3}=X^{24}+X^{57},&w_{4}=X^{27}+X^{45},\\[4.0pt] w_{5}=X^{34}+X^{67},&w_{6}=X^{37}+X^{46},&w_{7}=X^{25}-X^{47},&w_{8}=X^{36}-X^{47}.\end{array}$ (3.3)
For $\mathop{\mathrm{Re}}\Omega_{0}$ and $\mathop{\mathrm{Im}}\Omega_{0}$ given in Lemma 2.1, one verifies that $w_{i}\wedge\real\Omega_{0}=w_{i}\wedge\mathop{\mathrm{Im}}\Omega_{0}=0$ for each $i=1,\cdots,8$ in (3.3). Hence, if $A$ is a SDCI, then its curvature $F_{A}$ is locally generated by $w_{i}$’s in (3.3). Thus, for the $\mathrm{G}_{2}$-structure (cf. Corollary 2.8):
$\varphi=-\eta\wedge d\eta+\left(\frac{2}{z}\right)^{\frac{3}{4}}a\mathop{\mathrm{Re}}(\Omega_{0})+\left(\frac{2}{z}\right)^{\frac{3}{4}}b\mathop{\mathrm{Im}}(\Omega_{0}),$
we have
$\ast(\varphi\wedge F_{A})$ $=-\ast(\eta\wedge d\eta\wedge F_{A})+\left(\frac{2}{z}\right)^{\frac{3}{4}}a\ast(\mathop{\mathrm{Re}}(\Omega_{0})\wedge F_{A})+\left(\frac{2}{z}\right)^{\frac{3}{4}}b\ast(\mathop{\mathrm{Im}}(\Omega_{0})\wedge F_{A})$
$=-\ast(\eta\wedge d\eta\wedge F_{A})=-F_{A}.$
Remark 3.2.
We can generalise Lemma 3.1 in the following sense. Observe that from the decomposition in (3.2), we have the ${\rm SU}(3)$-invariant decomposition
$\Omega^{2}(H)=\Omega^{2}_{6}\oplus\Omega^{2}_{1}\oplus\Omega_{8}^{2},$ (3.4)
where $\Omega^{2}_{8}\simeq\mathfrak{su}(3)$ and $\Omega^{2}_{1}=\Omega^{0}(M)\otimes\omega$. On the other hand, we know that the space of $2$-forms on $H^{\mathbb{C}}=H\otimes_{\mathbb{R}}\mathbb{C}$ decomposes into $\operatorname{Gl}(3,\mathbb{C})$-invariant submodules
$\Omega^{2}(H^{\mathbb{C}})=\Omega^{2,0}(H^{\mathbb{C}})\oplus\Omega^{0,2}(H^{\mathbb{C}})\oplus\Omega^{1,1}(H^{\mathbb{C}}).$
By Schur’s Lemma, we have (see [falcitelli1994]*Equation (1.2)
$\Omega^{2,0}\oplus\Omega^{0,2}\simeq\Omega^{2}_{6}\otimes_{\mathbb{R}}\mathbb{C}\quad\text{and}\quad\Omega^{1,1}\simeq(\Omega^{2}_{1}\oplus\Omega^{2}_{8})\otimes_{\mathbb{R}}\mathbb{C}.$ (3.5)
On the other hand, we have the $\mathrm{G}_{2}$-invariant decomposition $\Omega^{2}(TM)=\Omega^{2}_{7}\oplus\Omega^{2}_{14},$ where $\Omega^{2}_{14}\simeq{\mathfrak{g}}_{2}$. Since ${\rm SU}(3)\subset\mathrm{G}_{2}$ we have that $\Omega_{8}^{2}\subset\Omega_{14}^{2}$ and $\Omega^{2}_{1}\subset\Omega^{2}_{7}$. It is easy to see that $\Omega^{2}_{V}$ has elements in both $\Omega_{7}^{2}$ and $\Omega_{14}^{2}$, hence the same is true for $\Omega^{2}_{6}$.
Lemma 3.3.
Let $E\to V^{5,2}$ be a Sasakian holomorphic vector bundle [Boyer2008, Section 3], with the Sasakian structure $(\eta,\xi,g,\Phi)$ given in Theorem 2.2 and $A\in\mathcal{A}(E)$ the Chern connection. If $A\in\mathcal{A}(E)$ is a $\mathrm{G}_{2}$-instanton then $A$ satisfies the equation (1.2), i.e. it is a $\lambda$-contact instanton .
Proof.
Consider $\varphi$ given in Corollary 2.8 and set $dz^{k}=X^{k}+\sqrt{-1}X^{k+3}$ for $k=2,3,4$. From Lemma 2.1 and (2.6), we know that $\Omega_{0}=dz^{2}\wedge dz^{3}\wedge dz^{4}\in\Lambda^{3,0}H^{\ast}_{\mathbb{C}}$, where $H:=\ker(\eta)$. Now, if $A\in\mathcal{A}(E)$ is the Chern connection, then $F_{A}$ is of $(1,1)$-type [biswas2010vector, p.552]. By taking into consideration the bi-degree of $\Omega_{0}$, we conclude that $F_{A}\wedge\Omega_{0}=0$ and $F_{A}\wedge\overline{\Omega_{0}}=0$, hence both $F_{A}\wedge\mathop{\mathrm{Im}}{\Omega_{0}}$ and $F_{A}\wedge\mathop{\mathrm{Re}}{\Omega_{0}}$ vanishes, then
$F_{A}\wedge\varphi$ $=F_{A}\wedge\left(\eta\wedge d\eta+\left(\frac{2}{z}\right)^{\frac{3}{4}}a\mathop{\mathrm{Re}}\Omega_{0}+\left(\frac{2}{z}\right)^{\frac{3}{4}}b\mathop{\mathrm{Im}}\Omega_{0}\right)$
$=F_{A}\wedge\eta\wedge d\eta.$
Consequently, solutions to the $\mathrm{G}_{2}$-instanton equation are solutions to the contact instanton equation. ∎
Consider the Kähler manifold $G_{2}(\mathbb{R}^{5})\cong V^{5,2}/S^{1}$, we fix a holomorphic bundle $(\mathcal{E},\bar{\partial})$ on $G_{5,2}$, the pullback $\mathcal{E}^{\ast}:=\pi^{\ast}\mathcal{E}\to V^{5,2}$ carries a structure of Sasakian holomorphic bundle
[…]
where $\pi\colon V^{5,2}\to G_{2}(\mathbb{R}^{5})$ is a Sasakian circle fibration endowed with the $\mathrm{G}_{2}$-structure (2.11). Finally, in [portilla2023, Lemma 2.4] was proved that SDCI are precisely connections whose curvature is of type $(1,1)$ and orthogonal to $\omega$, (cf. the splitting in (3.2) and Remark 3.2). Thus, from Lemmas 3.1 and 3.3, we obtain the following result:
Corollary 3.4.
The vector bundle $\mathcal{E}^{\ast}$ is a holomorphic Sasakian bundle, and a Chern connection $A$ on $\mathcal{E}$ is HYM if, and only if, $\pi^{\ast}A$ is a $\mathrm{G}_{2}$-instanton on $\mathcal{E}^{*}$.
We recall that in Lemma 3.1, Lemma 3.3 and Corollary 3.4 the key argument follow just in presence of a SU$(n)$-structure. These have analogous results in [portilla2023, Proposition 3.11] and [calvo2020gauge, Lemma 21], the difference in our case is that we do not assume $M$ is a contact Calabi-Yau but only endowed with a transversal ${\rm SU}(3)$-structure.
4 Homogeneous principal bundles on $V^{5,2}$
First, we recall some notation and results about homogeneous principal bundle, following [cap2009]*Section 1.4. Let $H\subset G$ be a closed subgroup of the Lie group $G$, we denote by $M=G/H$ the homogeneous space, $L\colon G\times M\to M$ the canonical left action and $p\colon G\to M$ an $H$–principal bundle given by the canonical projection.
Definition 4.1.
A homogeneous principal bundle over $M$ is a locally trivial principal bundle $\pi\colon P\to M$ together with a left $G$–action $\widetilde{L}\colon G\times P\to P$, which lifts the action $L$ on $M$, i.e., $\pi(\widetilde{L}({\textsl{g}},y))=L({\textsl{g}},\pi(y))$, and for each ${\textsl{g}}\in G$, the bundle map $\widetilde{L}_{\textsl{g}}\colon P\to P$ is a homomorphism of principal bundles.
Homogeneous principal bundles are characterised in the following way:
Lemma 4.2.
[cap2009, Lemma 1.4.5] Let $G$ and $K$ be Lie groups and $H\subset G$ a closed Lie subgroup. Let $P\to G/H$ a homogeneous $K$-principal bundle, then there exist a smooth homomorphism $\phi\colon H\to K$ such that $P\cong G\times_{\phi}K$. The action of $H$ on $K$ is given by $h\cdot k=\phi(h)\cdot k$ for each $h\in H$ and $k\in K$. The bundles corresponding to two homomorphisms $\phi,\widetilde{\phi}\colon H\to K$ are isomorphic if and only if there exist $k\in K$ such that $k\cdot\phi=\widetilde{\phi}\cdot k$
For a homogeneous $K$-principal bundle $P\to G/H$, a connection $A\in\Omega^{1}(P,{\mathfrak{k}})$ is called invariant, if $\widetilde{L}_{\textsl{g}}^{*}A=A$ for each ${\textsl{g}}\in G$.
Lemma 4.3.
[cap2009, Theorem 1.4.5] Let $P=G\times_{\phi}K\to G/H$ be the homogeneous $K$-principal bundle induced by the homomorphism $\phi:H\to K$. Then, invariant principal connections on $P$ are in bijective correspondence with linear maps $\alpha\colon{\mathfrak{g}}\to{\mathfrak{k}}$ such that
  • (i)
    $\alpha|_{{\mathfrak{h}}}=\phi_{\ast}\colon{\mathfrak{h}}\to{\mathfrak{k}}$, the derivative of $\phi$.
  • (ii)
    $\alpha\circ\mathrm{Ad}(h)=\mathrm{Ad}(\phi(h))\circ\alpha$ for all $h\in H$.
4.1 Homogeneous ${\rm SO}(3)$-principal bundles
Consider $H=K={\rm SO}(3)$ and $G={\rm SO}(5)$. Since ${\rm SO}(3)$ is simple, up to conjugation there are two smooth homomorphisms $\phi_{i}\colon{\rm SO}(3)\to{\rm SO}(3),$ $i=0,1$. Namely, $\phi_{0}$ the trivial homomorphism and $\phi_{1}=1_{{\rm SO}(3)}$ the identity, according to Lemma 4.2, there exist precisely two [^3] Up to isomorphism of homogeneous bundles homogeneous ${\rm SO}(3)$-principal bundles $\pi_{i}\colon P_{i}=G\times_{\phi_{i}}{\rm SO}(3)\to V^{5,2}$, $i=0,1$.
Lemma 4.4.
There is a unique invariant ${\rm SO}(3)$–connection on $P_{0}=V^{5,2}\times{\rm SO}(3)$, given by
$\alpha=e^{1}\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10}).$ (4.1)
Furthermore, the curvature $F_{\alpha}$ of the connection $\alpha$ belongs to $\Omega^{2}_{1}(\mathfrak{so}(3))$ cf. (3.4) and Remark 3.2.
Proof.
Consider an arbitrary linear map $\alpha:\mathfrak{so}(5)\rightarrow\mathfrak{so}(3)$, in the basis (2.2) we write
$\alpha=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}.$
From Lemma 4.3, $\alpha$ is an invariant connection, if and only if, it satisfies $\alpha|_{\mathfrak{so}(3)}=0$ and $\alpha\circ\mathrm{Ad}({\textsl{g}})=\alpha$, for all ${\textsl{g}}\in{\rm SO}(3)$. The condition $\alpha|_{\mathfrak{so}(3)}=0$ implies that $\alpha$ is of the form $\alpha=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$ for $i,j,k\in\{1,\dots,7\}$. Now, we apply the second condition to ${\textsl{g}}_{k}=\begin{pmatrix}I_{2\times 2}&\\ 0&h_{k}\\ \end{pmatrix}\in{\rm SO}(5),$ we define some values of $h_{k}$,
$h_{0}=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&-1\end{pmatrix},h_{1}=\begin{pmatrix}0&1&0\\ -1&0&0\\ 0&0&1\end{pmatrix},h_{2}=\begin{pmatrix}0&0&1\\ 0&-1&0\\ 1&0&0\end{pmatrix},\quad\text{and}\quad h_{3}=\begin{pmatrix}0&0&1\\ 0&1&0\\ -1&0&0\end{pmatrix}$ (4.2)
First, considering $h_{0}$ and $h_{1}$, we obtain that $\alpha(e_{i})=a_{i}e_{8}+b_{i}e_{9}+c_{i}e_{10}$ and, for $i\neq 3$ and $i\neq 6$
$\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{i}))=\alpha({\textsl{g}}_{0}e_{i}{\textsl{g}}_{0}^{-1})$ $=a_{i}e_{8}+b_{i}e_{10}+c_{i}e_{9}$
$\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{i}))=\alpha({\textsl{g}}_{1}e_{i}{\textsl{g}}_{1}^{-1})$ $=-a_{i}e_{8}-b_{i}e_{10}-c_{i}e_{9}$
follows that $a_{i}=b_{i}=c_{i}=0$ for $i\in\{2,4,5,7\}$. Now, by considering $h_{2}$ and $h_{3},$ we obtain for $i=3,6$
$\begin{array}[]{cl}\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{i}))=&-\alpha(e_{i})=a_{i}e_{8}+b_{i}e_{10}+c_{i}e_{9}\\[5.0pt] \alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{i}))=&\alpha(e_{i})=-a_{i}e_{8}-b_{i}e_{10}-c_{i}e_{9}.\end{array}$
Consequently, an invariant connection $\alpha$ is of the form
$\alpha=e^{1}\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10}).$ (4.3)
The curvature of $\alpha$ is $F_{\alpha}=y_{2}^{-1}(X^{25}+X^{36}+X^{47})\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10})$. The second affirmation in the Lemma follows from (2.6) since
$F_{\alpha}=2\omega\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10}).$ (4.4)
Remark 4.5 (Homogeneous ${\rm U}(1)$-connections).
Similar to Lemma 4.4, since ${\rm SO}(3)$ is simple, all non-trivial homomorphism $\phi\colon{\rm SO}(3)\rightarrow K$ is either trivial or injective for all Lie group $K$. In particular for $K={\rm U}(1)$, from Lemma 4.2, there exists a unique [^4] Up to isomorphism of principal bundles homogeneous ${\rm U}(1)$-bundle $P$ corresponding to $\phi$, namely, the trivial bundle $P=V^{5,2}\times{\rm U}(1)$. Any invariant connection on $P$ is given by a linear map $\beta:\mathfrak{so}(5)\rightarrow{\mathfrak{u}}(1)$ satisfying
$(i)\quad\beta|_{\mathfrak{so}(3)}=0=\phi_{\ast}\colon\mathfrak{so}(3)\to\mathbb{R}\quad\text{and}\quad(ii)\quad\beta\circ\mathrm{Ad}(a)=\beta,\text{ for all }a\in{\rm SO}(3).$
By considering $e_{1}$ the generator of the Abelian Lie algebra ${\mathfrak{u}}(1)=\mathbb{R}$, the unique invariant connection is given by
$\alpha=a_{i}e^{i}\otimes e_{1}\quad\text{for}\quad i=1,\dots,10.$
From item $(i)$, since $\alpha|_{\mathfrak{so}(3)}=0$, $\alpha$ has to be of the form $\alpha=a_{i}e^{i}\otimes e_{1}$, $i=1,\cdots,7.$ On the other hand, considering item $(ii)$ for ${\textsl{g}}_{0},{\textsl{g}}_{1}$ given in (4.2), we use the same argument used in the proof of Lemma 4.4 to obtain $\alpha(e_{i})=a_{i}e_{1}$ and
$\alpha(Ad({\textsl{g}}_{0})(e_{i}))=\alpha({\textsl{g}}_{0}e_{i}{\textsl{g}}_{0}^{-1}))$ $=-a_{i}e_{1}$
$\alpha(Ad({\textsl{g}}_{1})(e_{i}))=\alpha({\textsl{g}}_{1}e_{i}{\textsl{g}}_{1}^{-1}))$ $=a_{i}e_{1}.$
for $i=2,\cdots,7$. Therefore, $\alpha$ is characterised by
$\beta=a_{1}e^{1}\otimes e_{1}$ (4.5)
and its curvature $2$-form is
$F_{\beta}=2(X^{25}+X^{36}+X^{47})\otimes e_{1}.$ (4.6)
Note that $F_{\beta}\in\Omega^{2}_{1}=\omega\otimes\Omega^{0}(M)$ (4.13).
Now, we consider the ${\rm SO}(3)$-principal bundle $P_{1}={\rm SO}(5)\times_{\phi_{1}}{\rm SO}(3)$ induced by the homomorphism identity $\phi_{1}\colon{\rm SO}(3)\to{\rm SO}(3)$. The projection of $P_{1}$ is given by
$\pi_{1}\colon[(u,s)]\in P_{1}\mapsto u\cdot{\rm SO}(3)\in V^{5,2},$ (4.7)
and the class is given by $[(u,s)]=\{(u\cdot r,r^{-1}s)\;|\;r\in{\rm SO}(3)\}.$ There exist a left ${\rm SO}(5)$-action on $P_{1}$ such that for each $v\in{\rm SO}(5)$ the bundle map $L_{v}\colon P_{1}\to P_{1}$ is $\pi_{1}$-equivariant i.e.
$\pi_{1}(L_{v}[(u,s)])=\pi_{1}([(vu,s)])=vu\cdot{\rm SO}(3)=L_{v}(\pi_{1}[(u,s)]).$
According to (2.2), we fix the basis $\{e_{8},e_{9},e_{10}\}$ for $\mathfrak{so}(3)\subset\mathfrak{so}(5)$ such that $[e_{i},e_{j}]=\varepsilon_{ijk}e_{k}$, where $\varepsilon_{ijk}$ is the sign of the permutation $(i\;j\;k)$.
Lemma 4.6.
There is a unique invariant ${\rm SO}(3)$–connection on $P_{1}=V^{5,2}\times_{\phi_{1}}{\rm SO}(3)$ given by
$\alpha=e^{8}\otimes e_{8}+e^{9}\otimes e_{9}+e^{10}\otimes e_{10}.$ (4.8)
Proof.
Consider $\alpha:\mathfrak{so}(5)\rightarrow\mathfrak{so}(3)$ in the basis (2.2) given by $\alpha=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$. From Lemma 4.3 (i), we have that $\alpha=\alpha_{0}+e^{8}\otimes e_{8}+e^{9}\otimes e_{9}+e^{10}\otimes e_{10},$ where $\alpha_{0}=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$, $i,j,k\in\{1,\dots,7\}$, we are going to check that $\alpha_{0}=0$. From Lemma 4.3 (ii), for any ${\textsl{g}}\in{\rm SO}(3)$ and $X\in\mathfrak{so}(5)$, we have ${\textsl{g}}\alpha(X){\textsl{g}}^{-1}=\alpha({\textsl{g}}X{\textsl{g}}^{-1})$. Taking $X=e_{i}$ for each $i\in\{1,\dots,7\}$, we denote
$\alpha(\mathrm{Ad}({\textsl{g}})(e_{i}))=\alpha({\textsl{g}}e_{i}{\textsl{g}}^{-1})$ (4.9)
and
$\mathrm{Ad}({\textsl{g}})(\alpha(e_{i}))=a_{i}{\textsl{g}}e_{8}{\textsl{g}}^{-1}+b_{i}{\textsl{g}}e_{9}{\textsl{g}}^{-1}+c_{i}{\textsl{g}}e_{10}{\textsl{g}}^{-1}.$ (4.10)
Now, consider ${\textsl{g}}_{k}$ and $h_{k}$ as in (4.2), equations (4.9) and (4.10) become
$\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{2}))=\alpha({\textsl{g}}_{0}e_{2}{\textsl{g}}_{0}^{-1})=\alpha(e_{3})=a_{3}e_{8}+b_{3}e_{9}+c_{3}e_{10},$
and
$\mathrm{Ad}({\textsl{g}}_{0})(\alpha(e_{2}))=-a_{2}e_{8}+b_{2}e_{10}-c_{2}e_{9}.$
These imply $-a_{2}=a_{3},-c_{2}=b_{3},c_{3}=b_{2}$. Analogously, for $h_{1}$, we have
$\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{2}))=\alpha(e_{3})=-a_{3}e_{8}-b_{3}e_{9}-c_{3}e_{10}$
and
$\mathrm{Ad}({\textsl{g}}_{1})(\alpha(e_{2}))=a_{2}e_{8}+b_{2}e_{10}-c_{2}e_{9},$
consequently, $-a_{2}=a_{3},c_{2}=b_{3},-b_{2}=c_{3}$, hence $c_{2}=b_{3}=c_{3}=b_{2}=0$. Similarly:
  • From $\mathrm{Ad}({\textsl{g}}_{0})(\alpha(e_{4}))=\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{4}))$, we obtain $c_{4}=b_{4}=0$,
  • From $\mathrm{Ad}({\textsl{g}}_{1})(\alpha(e_{7}))=\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{7}))$, we obtain $c_{7}=b_{7}=0$.
  • From $\mathrm{Ad}({\textsl{g}}_{0})(\alpha(e_{5}))=\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{5}))$ and $\mathrm{Ad}({\textsl{g}}_{1})(\alpha(e_{5}))=\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{5}))$, we obtain $c_{6}=b_{5}=c_{5}=b_{6}=0$.
Finally, applying the same argument for $h_{2}$ and $h_{3},$ we conclude that $a_{i}=0$ for $i=1,\cdots,7$, consequently $\alpha_{0}=0$, then $\alpha=e^{8}\otimes e_{8}+e^{9}\otimes e_{9}+e^{10}\otimes e_{10}$ and (4.8) defines a unique ${\rm SO}(5)$-invariant connection $\alpha\in\Omega^{1}(P,\mathfrak{so}(3))$. ∎
In the basis (2.8), the curvature of (4.8) is given by
$F_{\alpha}=\frac{1}{y_{2}}\left((X^{23}+X^{56})\otimes e_{8}+(X^{24}+X^{57})\otimes e_{9}+(X^{34}+X^{67})\otimes e_{10}\right).$ (4.11)
Notice that $F_{\alpha}:{\mathfrak{m}}\times{\mathfrak{m}}\rightarrow\mathfrak{so}(3)$ is a ${\rm SO}(3)$-equivariant map, i.e., $F_{\alpha}\circ\mathrm{Ad}({\textsl{g}})=\mathrm{Ad}({\textsl{g}})\circ F_{\alpha}$ for any ${\textsl{g}}\in{\rm SO}(3)$, thus (4.11) corresponds with the ${\rm SO}(5)$-invariant curvature $2$-form of the induced by the ${\rm SO}(5)$-invariant connection $\alpha\in\Omega^{1}(P,\mathfrak{so}(3))$ (cf. [cap2009, Proposition 1.4.6]).
4.2 Homogeneous $\mathrm{G}_{2}$ and self dual contact instantons (SDCI)
For the ${\rm SO}(5)$-invariant contact form $\eta=X^{1}$ on $V^{5,2}$ and the ${\rm SO}(5)$-invariant metric given by (2.7), we consider the equation
$\ast_{g}(F_{\beta}\wedge\eta\wedge\omega)=\pm F_{\beta},\quad\text{where}\quad\omega=\frac{1}{2}d\eta=(X^{25}+X^{36}+X^{47}).$ (4.12)
From Lemma 4.6, we obtain the following
Corollary 4.7.
Consider $V^{5,2}$ with the Sasakian structure given in Theorem 2.2. Then, the connection (4.8) is an invariant SDCI on the homogeneous ${\rm SO}(3)$–principal bundle $P_{1}\to V^{5,2}$.
Proof.
For any Sasakian manifold, the bundle of $2$–forms decomposes into [portilla2023, Section 2.1]
$\Omega^{2}=\Omega^{2}_{1}\oplus\Omega^{2}_{6}\oplus\Omega^{2}_{8}\oplus\Omega^{2}_{V},$ (4.13)
where (3.3) are elements of $\Omega^{2}_{8}=span\{w_{i}\}_{i=1}^{8}$ solving $\ast(\eta\wedge\omega\wedge w_{i})=w_{i}$. Comparing (4.11) with the values of $w_{i}$ in (3.3), we note that $F_{\alpha}$ in (4.11) is given by
$F_{\alpha}:=\frac{1}{y_{2}}(w_{1}\otimes e_{8}+w_{3}\otimes e_{9}+w_{5}\otimes e_{10}).$
The curvature tensor $F_{\alpha}$ satisfies $\ast(F_{\alpha}\wedge\eta\wedge\omega)=F_{\alpha}$. Hence, for any $y_{2}\in\mathbb{R}^{+}$ satisfying (2.22), the connection (4.8) is an ${\rm SO}(5)$-invariant SDCI. ∎
Corollary 4.8.
Consider $V^{5,2}$ endowed with the Sasakian structure given in Corollary 2.8. There exist an unique invariant ASDCI on the trivial homogeneous ${\rm SO}(3)$–principal bundle $P_{0}\to V^{5,2}$ (cf. Lemma 4.4).
Proof.
From Lemma 4.4, we only should note that the curvature (4.4) satisfy $\ast_{g}(F_{\beta}\wedge\eta\wedge\omega)=F_{\beta},$ if and only if $y_{1}=\frac{9}{2}$. ∎
A result analogous to Corollary 4.8 follows for $P\to V^{5,2}$ in Remark 4.5, since the curvature tensor in (4.6) satisfy $\ast_{g}(F_{\beta}\wedge\eta\wedge\omega)=F_{\beta},$ if and only if $y_{1}=\frac{9}{2}$.
Proposition 4.9.
Consider the connection (4.8) on $V^{5,2}$, with respect to the $\mathrm{G}_{2}$–structure given in (2.11). Then we gave
  • The ${\rm SO}(3)$–connection in (4.8) on the homogeneous principal bundle $P_{1}\to V^{5,2}$ [cf. Lemma 4.6] satisfies $F_{\alpha}\wedge\psi=0$ for any $y_{1}>0$, i.e., there exist an unique ${\rm SO}(5)$–invariant connection which is simultaneously $\mathrm{G}_{2}$ and SDCI, given by (4.11).
  • The ${\rm SO}(3)$–connection (4.1) on the trivial bundle $P_{0}\to V^{5,2}$ (cf. Lemma 4.4), whose curvature is given in (4.4) satisfies $F_{\alpha}\wedge\psi=0$, for $\psi$ given in (2.15), if and only if (4.1) is flat.
4.3 Stability of SDCI and $\mathrm{G}_{2}$-instantons
Now, we analyse the Yang-Mills condition for instanton (4.8), for a detailed study on stability of Yang-Mills connections we refer to [bourguignon1981stability]. On the space of connections $\mathcal{A}$, the global inner product $(\alpha,\beta)=\int_{M}\langle\alpha\wedge\ast\beta\rangle$ allow us define the Yang-Mills functional as the $L^{2}$-norm of the curvature:
$\mathcal{YM}\colon\nabla\in\mathcal{A}\to\|{\rm F}_{\nabla}\|^{2}_{L^{2}}=\int_{M}\operatorname{Tr}({\rm F}_{\nabla}\wedge\ast{\rm F}_{\nabla})=\int_{M}\|{\rm F}_{\nabla}\|^{2}\in\mathbb{R}.$ (4.14)
A critical point of (4.14) is called a Yang-Mills (YM) connection, i.e. $\nabla$ satisfies the so-called Yang-Mills equation $d^{\ast}_{\nabla}F_{\nabla}=0$. Taking the exterior derivative $d_{\nabla}$ in (1.2) and using the Bianchi identity, one note that a solution of the instanton equation (1.2) satisfy the YM-equation with torsion:
$d_{\nabla}(\ast F_{\nabla})=d\sigma\wedge F_{\nabla}.$
Note that neither $\sigma=\eta\wedge d\eta$ nor $\sigma=\varphi$ are closed, consequently, the torsion term in the above equation is no automatically zero, however, we have that SDCI and ASDCI are Yang-Mills connections [luis2024Weitzenbock, Section 5]. With respect to the $\mathrm{G}_{2}$-structure (2.23), we have the following:
Theorem 4.1.
Let $\alpha$ be the unique $\mathrm{G}_{2}$-instanton (4.8) on the homogeneous ${\rm SO}(3)$-principal bundle $P_{1}$. Then, $\alpha$ is a critical point for the Yang-Mills energy for any coclosed $\mathrm{G}_{2}$-structure (2.23).
Proof.
Using (2.17) and the fact that $\alpha$ is a $\mathrm{G}_{2}$-instanton, we have that the curvature (4.11) satisfies $d_{\alpha}^{*}F_{\alpha}=0$, if and only if
$2a^{3}-2a^{2}x-3ab^{2}+2aby+bxy=0,\quad\text{and}\quad 2b^{3}-2b^{2}y-3a^{2}b+2abx+axy=0,$
we see that both equations hold for (2.23), i.e. $x=a$ and $y=b$. Hence, we obtain the result. ∎
For the Sasakian case, we can prove that SDCI are local minimizers of the YM-functional (4.14). In order to prove this, note that the contact structure induces an orthogonal decomposition (cf. [portilla2023, Lemma 2.3]) of the curvature $2$-form ${\rm F}_{\nabla}$ into
${\rm F}_{H}=i_{\xi}(\eta\wedge{\rm F}_{\nabla})\quad\text{and}\quad{\rm F}_{V}=\eta\wedge i_{\xi}{\rm F}_{\nabla},$
where $F_{H}$ is the horizontal part and $F_{V}$ the vertical part of $F_{\nabla}$. Consequently, we have $\|{\rm F}_{\nabla}\|^{2}=\|{\rm F}_{H}\|^{2}+\|{\rm F}_{V}\|^{2}$ and denote
${\rm F}_{H}^{+}=\frac{1}{2}({\rm F}_{H}+\ast(\sigma\wedge{\rm F}_{H}))\quad\text{and}\quad{\rm F}_{H}^{-}=\frac{1}{2}({\rm F}_{H}-\ast(\sigma\wedge{\rm F}_{H})),$ (4.15)
According to the splitting (3.2), we have ${\rm F}_{H}={\rm F}_{8}+{\rm F}_{6}+{\rm F}_{1}\in\Omega_{8}^{2}({\mathfrak{g}}_{E})\oplus\Omega_{6}^{2}({\mathfrak{g}}_{E})\oplus\Omega_{1}^{2}({\mathfrak{g}}_{E})$, then we observe that ${\rm F}_{H}^{+}={\rm F}_{8}-\frac{1}{2}{\rm F}_{1}$ and ${\rm F}_{H}^{-}={\rm F}_{6}+\frac{3}{2}{\rm F}_{1}$, hence
$\|{\rm F}_{H}\|^{2}$ $=\langle{\rm F}_{H}^{+}+{\rm F}_{H}^{-},{\rm F}_{H}^{+}+{\rm F}_{H}^{-}\rangle=\|{\rm F}_{H}^{+}\|^{2}+\|{\rm F}_{H}^{-}\|^{2}+2\langle{\rm F}_{H}^{+},{\rm F}_{H}^{-}\rangle$
$=\|{\rm F}_{H}^{+}\|^{2}+\|{\rm F}_{H}^{-}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2},$
consequently, the YM-functional (4.14) splits as
$\mathcal{YM}(\nabla)=\|{\rm F}_{H}^{+}\|^{2}+\|{\rm F}_{H}^{-}\|^{2}+\|{\rm F}_{V}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2}.$ (4.16)
In the Sasakian context, the charge of the connection $\nabla$ is defined by
$\kappa(\nabla)=\displaystyle\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}^{2})\wedge\eta\wedge\omega.$ (4.17)
Notice that the charge (4.17) is determined by the horizontal part of $F_{\nabla}$ [portilla2023, Proposition 3.13]:
$\kappa(\nabla)=\|{\rm F}_{H}^{+}\|^{2}-\|{\rm F}_{H}^{-}\|^{2}.$ (4.18)
Using (4.16) and (4.18), we have that the YM-mills action is given by
$\mathcal{YM}(\nabla)$ $=\kappa(\nabla)+2\|{\rm F}_{H}^{-}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2}+\|{\rm F}_{V}\|^{2}$
$=-\kappa(\nabla)+2\|{\rm F}_{H}^{+}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2}+\|{\rm F}_{V}\|^{2},$
hence, we have
$\mathcal{YM}(\nabla)=\begin{cases}\kappa(\nabla),&\nabla\in\Omega^{2}_{8}({\mathfrak{g}}_{E})\\ -\kappa(\nabla),&\nabla\in\Omega^{2}_{6}({\mathfrak{g}}_{E})\\ 2\kappa(\nabla),&\nabla\in\Omega^{2}_{1}({\mathfrak{g}}_{E}).\end{cases}$ (4.19)
We recall that if $M$ is a contact manifold, a $k$-form $\alpha\in\Omega^{k}(M)$ is called transversal form if $i_{\xi}\alpha=0$, we denote the space of transversal forms by $\Omega^{k}_{H}(M)$. Consequently, the space of $k$-forms splits as
$\Omega^{k}(M)=\Omega^{k}_{H}(M)\oplus\eta\wedge\Omega^{k-1}_{H}(M).$ (4.20)
More generally, consider a Riemannian manifold $M$ endowed with a Riemannian foliation $\mathcal{F}\subset M$ and consider a principal $G$-bundle $\pi\colon P\to M$. $P$ is said to be a foliated principal bundle if there exists a $G$-invariant foliation $\widetilde{\mathcal{F}}$ on $P$ such that the tangent distribution $T\widetilde{\mathcal{F}}\subset TP$ projects isomorphically onto $T\mathcal{F}$ under $\pi_{\ast}\colon TP\to TM$. Accordingly, a vector bundle is called foliated if its associated principal bundle is foliated. A local section $s\colon U\subset M\to P$ is said basic section if $s_{\ast}\colon TU\to TP$ maps $T\mathcal{F}$ to $T\widetilde{\mathcal{F}}$. It is known that every foliated principal bundle admits local basic sections [molino1988riemannian, Proposition 2.7]. We denote the space of basic sections by $\Gamma(E)_{B}\subset\Gamma(E)$. In the context of contact manifolds, $\mathcal{F}$ is naturally given by the Reeb foliation, if $\nabla\in\Gamma(E)$ is a connection with $i_{\xi}F_{\nabla}=0$, then the horizontal lift of $\mathcal{F}$ with respect to $\nabla$ is an integrable distribution and gives $E$ a foliated structure for which $\nabla$ is basic connection [molino1988riemannian], consequently, we have a natural split of the bundle of $E$-valued $k$-forms analogous to (4.20).
On the underlying bundle $E$, considering only transverse deformations, i.e., $\nabla^{\prime}=\nabla+\alpha$, where $\alpha\in\Omega^{1}_{H}({\mathfrak{g}}_{E})$, in [portilla2023, Lemma 3.14] is shown that the charge vary by
$\kappa(\nabla)=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\alpha)\wedge\omega^{2}.$ (4.21)
The following result emulates [portilla2023, Lemma 3.14], the proof proceeds with the same reasoning, but we emphasise that, first, this result is not exclusive to the Stiefel manifold, but rather applies to any Sasakian manifold. And secondly, in [portilla2023] the result is announced for transverse deformations of the Chern connection [portilla2023, Corollary 3.15], but actually this result holds in a more general context as follow:
Lemma 4.10.
Let $E$ be a Sasakian ${\rm G}$-bundle over a compact closed 7-manifold. For any variation $\nabla^{\prime}=\nabla+\alpha\in\mathcal{A}(E)$, $\alpha\in\Omega^{1}({\mathfrak{g}}_{E})$, on the underlying complex vector bundle E, we have
$\kappa(\nabla)=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\ast(\omega\otimes s)),$ (4.22)
where $s\in\Gamma(E)_{B}$ is some transversal section of $E$.
Proof.
The proof of [portilla2023, Lemma 3.14] provide that for an arbitrary deformation $\nabla^{\prime}=\nabla+\alpha$, where $\alpha\in\Omega^{1}({\mathfrak{g}}_{E})=\Omega^{1}_{H}({\mathfrak{g}}_{E})\oplus\eta\wedge\Omega^{0}_{H}({\mathfrak{g}}_{E})$, the charge varies as
$\kappa(\nabla)$ $=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\alpha)\wedge\omega^{2}$ (4.23)
$\quad+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(d_{\nabla}\alpha\wedge\alpha)\wedge\omega^{2}+\frac{1}{3}\int_{M}\mathop{\mathrm{Tr}}\nolimits(\alpha\wedge\alpha\wedge\alpha)\wedge\omega^{2}.$
In particular, assuming a transversal deformation, (4.21) holds because the defect term in (4.23) vanishes, as the subintegral quantities in the last two terms belong to $\Omega^{7}_{H}({\mathfrak{g}}_{E})=0$. Hence, we can assume that $\alpha$ is a vertical $1$-form, i.e., $\alpha=\eta\otimes s\in\eta\wedge\Omega^{0}_{H}({\mathfrak{g}}_{E})$ for some transversal section $s$, therefore
$\kappa(\nabla)$ $=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge(\eta\otimes s))\wedge\omega^{2}+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(d_{\nabla}(\eta\otimes s)\wedge(\eta\otimes s))\wedge\omega^{2}$
$=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\eta\wedge\omega^{2}\otimes s)+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(\omega\otimes s-\eta\otimes\nabla s)\wedge(\eta\otimes s))\wedge\omega^{2}$
$=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\ast(\omega\otimes s))+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(\eta\wedge\omega^{3}\otimes[s,s]),$
hence, (4.22) follows. ∎
Corollary 4.11.
Let $E$ be a Sasakian ${\rm G}$-bundle over a compact closed 7-manifold. Any SDCI is a local minimizer of the YM-functional (4.14).
Proof.
From (4.19), note that the YM-functional assumes the value of the charge $\kappa(\nabla)$ (4.17). From (4.22)
$\kappa(\nabla)$ $=\kappa(\nabla^{\prime})+({\rm F}_{\nabla},\omega\otimes s)_{L^{2}}$
$=\kappa(\nabla^{\prime}),$
the last line follows since $\Omega^{2}_{8}({\mathfrak{g}}_{E})\perp\omega\otimes\Omega^{0}({\mathfrak{g}}_{E})$. Hence, if $\nabla$ is a SDCI the charge is locally constant. The same is true if $\nabla$ is a ASDCI, since $\Omega^{2}_{6}({\mathfrak{g}}_{E})\perp\omega\otimes\Omega^{0}({\mathfrak{g}}_{E})$. ∎

### v2

@@
 Abstract
-We study homogeneous instantons on the seven dimensional Stiefel manifold $V^{5,2}$ in the context of $\mathrm{G}_{2}$ and Sasakian geometry.
+We study homogeneous instantons on the seven–dimensional Stiefel manifold $V^{5,2}$ in the context of $\mathrm{G}_{2}$ and Sasakian geometry.
 According to the reductive decomposition of $V^{5,2}$, we provide an explicit description of all invariant $\mathrm{G}_{2}$ and Sasakian structures.
-In particular, we characterise the invariant $\mathrm{G}_{2}$-structures inducing a Sasakian metric, among which the well known nearly parallel $\mathrm{G}_{2}$-structure (Sasaki-Einstein) is included.
+In particular, we characterise the invariant $\mathrm{G}_{2}$–structures inducing a Sasakian metric, among which the well–known nearly parallel $\mathrm{G}_{2}$–structure (Sasaki–Einstein) is included.
 As a consequence, we classify the invariant connections on homogeneous principal bundles over $V^{5,2}$ with gauge group ${\rm U}(1)$ and ${\rm SO}(3)$, satisfying either the $\mathrm{G}_{2}$ or the Sasakian instanton condition.
-Finally, we analyse the Yang Mills condition for those invariant connections.
+In addition, we study infinitesimal deformations of $\mathrm{G}_{2}$–instantons on coclosed $\mathrm{G}_{2}$–manifolds using a spinorial approach.
+By means of a Weitzenböck–type formula with torsion, we obtain curvature obstructions to the existence of non–trivial infinitesimal deformations and prove rigidity results for certain homogeneous $\mathrm{G}_{2}$–instantons.
 1 Introduction
 The Stiefel manifold $V^{n,k}$ of orthonormal $k$-frames in the $n$-dimensional Euclidean space (i.e.
-$1\leq k\leq n$) admits a transitive ${\rm SO}(n)$-action with isotropy subgroup isomorphic to ${\rm SO}(n-k)$, thus, it can be written homogeneously as $V^{n,k}={\rm SO}(n)/{\rm SO}(n-k)$ and consequently, its family of ${\rm SO}(n)$-invariant Riemannian metrics are described in terms of the isotropy representation of $V^{n,k}$.
+$1\leq k\leq n$) admits a transitive ${\rm SO}(n)$-action with isotropy subgroup isomorphic to ${\rm SO}(n-k)$, thus, it can be written homogeneously as $V^{n,k}={\rm SO}(n)/{\rm SO}(n-k)$ and consequently, its family of ${\rm SO}(n)$-invariant Riemannian metrics is described in terms of the isotropy representation of $V^{n,k}$.
 For instance, for $k=1$ or $k=n$, the isotropy representation is irreducible (indeed $V^{n,1}=S^{n}$ or $V^{n,n}={\rm SO}(n)$), in those cases, there is a unique ${\rm SO}(n)$-invariant metric, which is Einstein [wolf1968].
-For $k=2$, the isotropy representation is reducible and the space of ${\rm SO}(n)$-invariant Riemannian metrics is described by a $4$-dimensional family, within there is a unique Einstein metric, which is inherited from the Grasmmanian $G_{2}(\mathbb{R}^{n})$ (e.g.
+For $k=2$, the isotropy representation is reducible and the space of ${\rm SO}(n)$-invariant Riemannian metrics is described by a $4$-dimensional family, within which there exists a unique Einstein metric, which is inherited from the Grassmannian $G_{2}(\mathbb{R}^{n})$ (e.g.
 [Arvanitoyeorgos1996, kerr1998new]).
-In addition, $V^{n,2}$ fits as an instance of the Boothby-Wang construction, as a total space of a $S^{1}$-principal bundle [boothby1958]*Theorem 3, moreover, it has a regular contact structure whose associated Reeb vector field generates a $1$-parameter subgroup of right translations of $V^{n,2}$.
-A contact structure [^1] The reader unfamiliar with Sasakian geometry can find a complete study on this subject in [Boyer2008] on a $n$-dimensional manifold $M$ is a 1-form satisfying $\eta\wedge(d\eta)^{n}\neq 0$, in this case there exist a unique vector $\xi\in\mathscr{X}(M)$, called Reeb vector field, such that $\eta(\xi)=1$.
-The contact structure is called an almost contact structure if there exist an endomorphism $\Phi\in{\mathrm{End}}(TM)$ which satisfy $\Phi^{2}=-{\rm I}_{TM}+\eta\otimes\xi.$ A Riemannian metric $g$ is called compatible with the contact structure, if $g$ satisfies $g(\Phi X,\Phi Y)=g(X,Y)-\eta(X)\eta(Y),$ in this case, the contact structure is called an almost contact metric structure.
+In addition, $V^{n,2}$ fits as an instance of the Boothby-Wang construction, as the total space of a $S^{1}$-principal bundle [boothby1958]*Theorem 3, moreover, it has a regular contact structure whose associated Reeb vector field generates a one-parameter subgroup of right translations of $V^{n,2}$.
+A contact structure [^1] The reader unfamiliar with Sasakian geometry can find a complete study on this subject in [Boyer2008] on a $(2n+1)$-dimensional manifold $M$ is a 1-form satisfying $\eta\wedge(d\eta)^{n}\neq 0$, in this case there exists a unique vector $\xi\in\mathscr{X}(M)$, called Reeb vector field, such that $\eta(\xi)=1$.
+The contact structure is called an almost contact structure if there exists an endomorphism $\Phi\in{\mathrm{End}}(TM)$ which satisfy $\Phi^{2}=-{\rm I}_{TM}+\eta\otimes\xi.$ A Riemannian metric $g$ is called compatible with the contact structure, if $g$ satisfies $g(\Phi X,\Phi Y)=g(X,Y)-\eta(X)\eta(Y),$ in this case, the contact structure is called an almost contact metric structure.
 In addition, if $\mathcal{S}=(\eta,\xi,g,\Phi)$ satisfies $2g(X,\Phi(Y))=d\eta(X,Y)$, the quadruple $\mathcal{S}$ is called a contact metric structure.
 Furthermore, if the Nijenhuis tensor $N_{\Phi}$ induced by $\Phi$ satisfies $N_{\Phi}=-d\eta\otimes\xi$, then $\mathcal{S}$ is called a Sasakian structure.
-In this paper, we consider the $7$–dimensional Stiefel manifold $V^{5,2}$ endowed with a one-parameter family of ${\rm SO}(5)$-invariant Sasakian structures $\mathcal{S}=(\eta,\xi,g,\Phi)$.
-For instance, homogeneous contact/Sasakian structures have been study in [andrada2009, correa2019] and [perrone1998homogeneous].
-On the other hand, a differential $3$-form $\varphi$ on an oriented $7$–dimensional manifold $M$ is called a $\mathrm{G}_{2}$–structure if at each $p\in M$, there is a basis $e_{1},\dots,e_{7}$ of $T_{p}M$ such that
+In this paper, we consider the 77–dimensional Stiefel manifold $V^{5,2}$ endowed with a one-parameter family of ${\rm SO}(5)$-invariant Sasakian structures $\mathcal{S}=(\eta,\xi,g,\Phi)$.
+For instance, homogeneous contact/Sasakian structures have been studied in [andrada2009, correa2019] and [perrone1998homogeneous].
+On the other hand, a differential $3$-form $\varphi$ on an oriented 77–dimensional manifold $M$ is called a $\mathrm{G}_{2}$–structure if at each $p\in M$, there is a basis $e_{1},\dots,e_{7}$ of $T_{p}M$ such that
 $\varphi_{p}=e^{125}+e^{136}+e^{147}+e^{234}-e^{267}+e^{357}-e^{456},$ (1.1)
@@
 In particular, $\varphi$ determines a Riemannian metric $g_{\varphi}$, we study the family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$, such that the induced metric $g_{\varphi}$ is Sasakian.
-With respect to these structures, we study the theoretical Gauge instanton equations associated with the underlying Sasakian/$\mathrm{G}_{2}$-geometry of $V^{5,2}$.
-Namely, on a $G$–principal bundle $P\to M^{7}$ with $\sigma\in\Omega^{3}(M)$, the $1$-form connection $A\in\mathcal{A}(P)$ is called an $\lambda$-instanton, if it satisfies (e.g.
+With respect to these structures, we study the theoretical gauge instanton equations associated with the underlying Sasakian and $\mathrm{G}_{2}$-geometry of $V^{5,2}$.
+Namely, on a GG–principal bundle $P\to M^{7}$ with $\sigma\in\Omega^{3}(M)$, the $1$-form connection $A\in\mathcal{A}(P)$ is called an $\lambda$-instanton, if it satisfies (e.g.
 [donaldson1998gauge, Tian2000])
@@
   1. 1.
-    A Self-dual contact Instantons (SDCI) [portilla2023] is a solutions of (1.2) with $\lambda=1$ and $\sigma:=\eta\wedge\omega$ naturally defined by a contact structure of $(M,\eta)$, where $\omega:=\frac{1}{2}d\eta$ is called the fundamental $2$-form..
+    A Self-dual contact instantons (SDCI) [portilla2023] is a solution of (1.2) with $\lambda=1$ and $\sigma:=\eta\wedge\omega$ naturally defined by a contact structure of $(M,\eta)$, where $\omega:=\frac{1}{2}d\eta$ is called the fundamental $2$-form.
   2. 2.
-    A $\mathrm{G}_{2}$–instantons [corrigan1983first, donaldson1998gauge] is a solutions of (1.2) with $\lambda=-1$ and $\sigma=\varphi$ a coclosed (i.e.
+    A $\mathrm{G}_{2}$–instantons [corrigan1983first, donaldson1998gauge] is a solution of (1.2) with $\lambda=-1$ and $\sigma=\varphi$ a coclosed (i.e.
 $d\ast\varphi=0$) $\mathrm{G}_{2}$–structure on $M$.
+Equivalently, a connection $A$ is a $\mathfrak{g}$}–instanton if and only if
+    $F_{A}\wedge\ast\varphi=0$ (1.3)
 Furthermore, for $E\to(M^{7},\mathcal{S})$ a Sasakian holomorphic bundle [biswas2010vector], the usual notion of Hermitian Yang-Mills (HYM) connection on a Kähler manifolds is extended to odd dimension in the natural fashion to the so-called transverse Hermitian Yang-Mills connections (tHYM) (see (3.1)).
-The concept of Chern connections also extend naturally to a Sasakian version, as a connection mutually compatible with the transverse holomorphic structure and the Hermitian metric.
+The concept of Chern connections also extends naturally to a Sasakian version, as a connection mutually compatible with the transverse holomorphic structure and the Hermitian metric.
 A noteworthy result is that if $E\to M$ is a holomorphic Sasakian bundle on a contact Calabi-Yau manifold $M$, along compatible connections the notion of SDCI, $\mathrm{G}_{2}$-instantons and tHYM coincide.
 cf.
-[portilla2023, Teorem 1.1] and [calvo2020gauge, Lemma  21].
+[portilla2023, Theorem 1.1] and [calvo2020gauge, Lemma  21].
 In [donaldson1998gauge, Donaldson2011], Donaldson, Segal and Thomas suggested that it may be possible to use $\mathrm{G}_{2}$-instantons to construct an enumerative invariant of $\mathrm{G}_{2}$-manifolds.
 For instance, when $(M^{7},\varphi)$ has a torsion-free $\mathrm{G}_{2}$-structure $\varphi$ (i.e.
-$d\varphi=0$ and $d\ast\varphi=0$ [fernandez1982riemannian]), solutions to the $\mathrm{G}_{2}$-instanton equation have been studied in [sa2015g2, walpuski2015g_2, clarke2014instantons, lotay20182].
-However, the construction of metrics with holonomy $\mathrm{G}_{2}$ has proven to be a challenging task, thus, the torsion-free condition has been relaxed to the class of coclosed $\mathrm{G}_{2}$-structures [ball2019gauge, clarke2023], which has a notably application in theoretical physics to the Killing spinor equations in supergravity (cf.
+$d\varphi=0$ and $d\ast\varphi=0$ [fernandez1982riemannian]), solutions to the $\mathrm{G}_{2}$-instanton equation have been studied in [sa2015g2, walpuski2015g_2, clarke2014instantons, lotay20182, singhal2022].
+However, the construction of metrics with holonomy $\mathrm{G}_{2}$ has proven to be a challenging task, thus, the torsion-free condition has been relaxed to the class of coclosed $\mathrm{G}_{2}$-structures [ball2019gauge, clarke2023], which has a notable application in theoretical physics to the Killing spinor equations in supergravity (cf.
 [Ivanov2005, Ivanov10, delaOssaG21, Lotay2022, Fernandez2011]).
@@
 $a^{2}+by>0,\quad b^{2}+ax>0\quad z\neq 0\quad\text{and}\quad xy=ab.$
-In particular, it defines a family of (non-purely) coclosed $\mathrm{G}_{2}$-structures, within the nearly parallel $\mathrm{G}_{2}$-structures arises as an $S^{1}$-family for $x=a$, $y=b$ and $a^{2}+b^{2}=z^{2}=\frac{27}{512}$ (cf.
+In particular, it defines a family of (non-purely) coclosed $\mathrm{G}_{2}$-structures, within the nearly parallel $\mathrm{G}_{2}$-structures arises as an $S^{1}$–family for $x=a$, $y=b$ and $a^{2}+b^{2}=z^{2}=\frac{27}{512}$ (cf.
 Theorem 2.1).
@@
 Conversely, for a holomorphic Sasakian bundle $E\to V^{5,2}$, if the Chern connection is a $\mathrm{G}_{2}$-instanton, then it is a SCDI.
-Finally, we prove that HYM connections on a holomorphic bundle $E\to G_{2}(\mathbb{R}^{5})$ on the Grassmanian, pulling back to $\mathrm{G}_{2}$-instantons on $\pi^{\ast}E\to V^{5,2}$ (cf.
+Finally, we prove that HYM connections on a holomorphic bundle $E\to G_{2}(\mathbb{R}^{5})$ on the Grassmannian, pulling back to $\mathrm{G}_{2}$-instantons on $\pi^{\ast}E\to V^{5,2}$ (cf.
 Corollary 3.4).
-We recall here that the key argument follow just in presence of a SU$(3)$-structure.
-Finally, in Section 4 we provide examples of homogeneous SCDI and $\mathrm{G}_{2}$-instantons over $V^{5,2}$, with gauge group ${\rm SO}(3)$ and ${\rm U}(1)$.
+We point out that the key argument can be carried out under the sole assumption of an $\mathrm{SU}(3)$-structure, thereby improving the results mentioned above.
+In Section 4, we provide examples of homogeneous SDCI and $\mathrm{G}_{2}$-instantons over $V^{5,2}$, with gauge group ${\rm SO}(3)$ and ${\rm U}(1)$.
 In particular, we prove that the ${\rm SO}(3)$-invariant $\mathrm{G}_{2}$-instanton (4.9) is a Yang-Mills connection on $(V^{5,2},\varphi)$, for any ${\rm SO}(5)$-invariant coclosed $\mathrm{G}_{2}$-structure (cf.
-Theorem 4.1), moreover, it describes a local minimizer of the Yang-Mills functional (cf.
-Corollary 4.11).
+Theorem 4.1).
+Finally, in Section 5 we extend the spinorial framework proposed by R.
+Singhal [singhal2022] to coclosed $\mathrm{G}_{2}$–structures, and as a concrete application in Corollary 5.7, we prove rigidity for a family of homogeneous $\mathrm{G}_{2}$–instantons on the Stiefel manifold $V^{5,2}$ for a non-empty range of the metric parameter $y_{2}$.
 Outline: In Section 2, we provide some preliminaries about $V^{5,2}$ as a homogeneous space, with the aim of characterising all the ${\rm SO}(5)$-invariant coclosed $\mathrm{G}_{2}$-structures and the associated Sasakian metrics.
-In Section 3, we explore the interaction between SDCI and $\mathrm{G}_{2}$–instantons on $V^{5,2}$, in particular, Lemmata 3.1, 3.3 and Corollary 3.4 indicate conditions such that contact instantons and $\mathrm{G}_{2}$-instantons coincides.
-Finally, in Section 4 we consider homogeneous principal bundles on $V^{5,2}$ and we construct examples of invariant instantons on ${\rm SO}(3)$, $U(1)$–homogeneous bundles.
-In particular, in Section 4.1 we characterise all invariant ${\rm SO}(3)$–connection and we show that these satisfy the $\mathrm{G}_{2}$ and SDCI condition for appropriate parameters and we analyse its stability as Yang-Mills connections.
-Acknowledgments and funding: AM was funded by the São Paulo Research Foundation (Fapesp) [2021/08026-5] and [2023/13780-6].
-LP was funded by Centre Henri Lebesgue (CHL) under the grant ANR-11-LABX-0020-01, he also would like to thanks to LMBA at Université de Bretagne Occidentale UBO for the hospitality.
+In Section 3, we explore the interaction between SDCI and $\mathrm{G}_{2}$–instantons on $V^{5,2}$, in particular, Lemmata 3.1, 3.3 and Corollary 3.4 indicate conditions such that contact instantons and $\mathrm{G}_{2}$-instantons coincide.
+Finally, in Section 4 we consider homogeneous principal bundles on $V^{5,2}$ and we construct examples of invariant instantons on ${\rm SO}(3)$, U(1)U(1)–homogeneous bundles.
+In particular, in Section 4.1 we characterise all invariant ${\rm SO}(3)$–connection and we show that these satisfy the $\mathrm{G}_{2}$ and SDCI condition for appropriate parameters.
+Finally, in Section 5 we study the deformation theory for a particular example of a $\mathrm{G}_{2}$-instanton, specifically in Corollary 5.7 we prove that it is rigid for a non-empty interval of the metric parameter $y_{2}$.
+Acknowledgments and funding: AM was funded by the São Paulo Research Foundation [2021/08026-5] and [2023/13780-6] (Fapesp).
+LP was funded by Centre Henri Lebesgue (CHL) under the grant ANR-11-LABX-0020-01.
+He also thanks the LMBA at Université de Bretagne Occidentale (UBO) for its hospitality
 2 Homogeneous $\mathrm{G}_{2}$ and contact metric structures on $V^{5,2}$
 In this section, we provide an explicit description of a $\mathrm{Ad}({\rm SO}(3))$-invariant $\mathrm{G}_{2}$-structure and the induced inner product on $V^{5,2}$.
-As well as, we establish the notation that we will use throughout this work.
+We also establish the notation that we will use throughout this work.
 2.1 $\mathrm{G}_{2}$-structures on $V^{5,2}$
-In order to describe the space of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structure on $V^{5,2}={\rm SO}(5)/{\rm SO}(3)$ [^2] This subject is brought up in [friedrich1997nearly]., we fix a reductive decomposition of $V^{5,2}$ as a homogeneous space.
+In order to describe the space of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}={\rm SO}(5)/{\rm SO}(3)$ [^2] This subject is discussed in depth in [friedrich1997nearly]., we fix a reductive decomposition of $V^{5,2}$ as a homogeneous space.
 Consider the Lie algebra of ${\rm SO}(5)$
 $\mathfrak{so}(5):=\{A\in\mathfrak{gl}(\mathbb{R}^{5})\;|\;A+A^{T}=0\},$ (2.1)
-and denote by $E_{ij}$, the $5\times 5$ matrix with $1$ at the $(ij)$-entry and 00 elsewhere, and fix a basis for $\mathfrak{so}(5)$:
-$\begin{split}e_{1}=&E_{12}-E_{21},\quad e_{4}=E_{15}-E_{51},\quad e_{7}=E_{25}-E_{52},\quad e_{10}=E_{54}-E_{45},\\ e_{2}=&E_{13}-E_{31},\quad e_{5}=E_{23}-E_{32},\quad e_{8}=E_{34}-E_{43},\quad\\ e_{3}=&E_{14}-E_{41},\quad e_{6}=E_{24}-E_{42},\quad e_{9}=E_{35}-E_{53}.\end{split}$ (2.2)
-We consider the Lie algebra $\mathfrak{so}(3)$ embedded in $\mathfrak{so}(5)$ and it is spanned by $\{e_{8},e_{9},e_{10}\}$.
-The Lie bracket of $\mathfrak{so}(5)$ is given by the commutator describe in Table 1.
-[,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] $e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $e_{6}$ $e_{7}$ $e_{8}$ $e_{9}$ $e_{10}$
-$e_{1}$ 00 $-e_{5}$ $-e_{6}$ $-e_{7}$ $e_{2}$ $e_{3}$ $e_{4}$ 00 00 00
-$e_{2}$ $e_{5}$ 00 $-e_{8}$ $-e_{9}$ $-e_{1}$ 00 00 $e_{3}$ $e_{4}$ 00
-$e_{3}$ $e_{6}$ $e_{8}$ 00 $-e_{10}$ 00 $-e_{1}$ 00 $-e_{2}$ 00 $e_{4}$
-$e_{4}$ $e_{7}$ $e_{9}$ $e_{10}$ 00 00 00 $-e_{1}$ 00 $-e_{2}$ $-e_{3}$
-$e_{5}$ $-e_{2}$ $e_{1}$ 00 00 00 $-e_{8}$ $-e_{9}$ $e_{6}$ $e_{7}$ 00
-$e_{6}$ $-e_{3}$ 00 $e_{1}$ 00 $e_{8}$ 00 $-e_{10}$ $-e_{5}$ 00 $e_{7}$
-$e_{7}$ $-e_{4}$ 00 00 $e_{1}$ $e_{9}$ $e_{10}$ 00 00 $-e_{5}$ $-e_{6}$
-$e_{8}$ 00 $-e_{3}$ $e_{2}$ 00 $-e_{6}$ $e_{5}$ 00 00 $-e_{10}$ $e_{9}$
-$e_{9}$ 00 $-e_{4}$ 00 $e_{2}$ $-e_{7}$ 00 $e_{5}$ $e_{10}$ 00 $-e_{8}$
-$e_{10}$ 00 00 $-e_{4}$ $e_{3}$ 00 $-e_{7}$ $e_{6}$ $-e_{9}$ $e_{8}$ 00
+Let $E_{ij}$ denote the elementary $5\times 5$ matrix with a $1$ in the $(i,j)$-entry and zeros elsewhere. We fix the following basis of the Lie algebra $\mathfrak{so}(5)$:
+$\begin{split}e_{1}=&E_{12}-E_{21},\quad e_{4}=E_{15}-E_{51},\quad e_{7}=E_{25}-E_{52},\quad e_{10}=E_{54}-E_{45},\\[3.00003pt] e_{2}=&E_{13}-E_{31},\quad e_{5}=E_{23}-E_{32},\quad e_{8}=E_{34}-E_{43},\quad\\[3.00003pt] e_{3}=&E_{14}-E_{41},\quad e_{6}=E_{24}-E_{42},\quad e_{9}=E_{35}-E_{53}.\end{split}$ (2.2)
+We consider the Lie algebra $\mathfrak{so}(3)$ embedded in $\mathfrak{so}(5)$ which is spanned by $\{e_{8},e_{9},e_{10}\}$.
+The Lie bracket of $\mathfrak{so}(5)$ is given by the commutator described in Table 1.
+$\cdot$$\cdot$] $e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $e_{6}$ $e_{7}$ $e_{8}$ $e_{9}$ $e_{10}$
+$e_{1}$ 0 $e_{5}$ $e_{6}$ $e_{7}$ $e_{2}$ $e_{3}$ $e_{4}$ 0 0 0
+$e_{2}$ $e_{5}$ 0 $e_{8}$ $e_{9}$ $e_{1}$ 0 0 $e_{3}$ $e_{4}$ 0
+$e_{3}$ $e_{6}$ $e_{8}$ 0 $e_{10}$ 0 $e_{1}$ 0 $e_{2}$ 0 $e_{4}$
+$e_{4}$ $e_{7}$ $e_{9}$ $e_{10}$ 0 0 0 $e_{1}$ 0 $e_{2}$ $e_{3}$
+$e_{5}$ $e_{2}$ $e_{1}$ 0 0 0 $e_{8}$ $e_{9}$ $e_{6}$ $e_{7}$ 0
+$e_{6}$ $e_{3}$ 0 $e_{1}$ 0 $e_{8}$ 0 $e_{10}$ $e_{5}$ 0 $e_{7}$
+$e_{7}$ $e_{4}$ 0 0 $e_{1}$ $e_{9}$ $e_{10}$ 0 0 $e_{5}$ $e_{6}$
+$e_{8}$ 0 $e_{3}$ $e_{2}$ 0 $e_{6}$ $e_{5}$ 0 0 $e_{10}$ $e_{9}$
+$e_{9}$ 0 $e_{4}$ 0 $e_{2}$ $e_{7}$ 0 $e_{5}$ $e_{10}$ 0 $e_{8}$
+$e_{10}$ 0 0 $e_{4}$ $e_{3}$ 0 $e_{7}$ $e_{6}$ $e_{9}$ $e_{8}$ 0
 Table 1: Bracket of $\mathfrak{so}(5)$ with respect to the basis (2.2)
@@
 $\mathrm{ad}(e_{9})_{\mathfrak{m}}$ $=\mathrm{diag}(0,A_{2},A_{2}),$
-$\mathrm{ad}(e_{10})_{\mathfrak{m}}$ $=\mathrm{diag}(0,A_{3}A_{3}),$
+$\mathrm{ad}(e_{10})_{\mathfrak{m}}$ $=\mathrm{diag}(0,A_{3},A_{3}),$
 where $A_{1}=\begin{pmatrix}0&-1&0\\ 1&0&0\\ 0&0&0\end{pmatrix}$,   $A_{2}=\begin{pmatrix}0&0&-1\\ 0&0&0\\ 1&0&0\end{pmatrix}$ and $A_{3}=\begin{pmatrix}0&0&0\\ 0&0&-1\\ 0&1&0\end{pmatrix}$ are the canonical basis of $\mathfrak{so}(3)$.
-Denote by ${\mathfrak{m}}_{1}=\operatorname{{\rm Span}}\{e_{1}\},$ ${\mathfrak{m}}_{2}:=\operatorname{{\rm Span}}\{e_{2},e_{3},e_{4}\}$ and ${\mathfrak{m}}_{3}:=\operatorname{{\rm Span}}\{e_{5},e_{6},e_{7}\}.$ By using the Lie bracket in Table 1 is straightforward to verify that
+Denote by ${\mathfrak{m}}_{1}=\operatorname{{\rm Span}}\{e_{1}\},$ ${\mathfrak{m}}_{2}:=\operatorname{{\rm Span}}\{e_{2},e_{3},e_{4}\}$ and ${\mathfrak{m}}_{3}:=\operatorname{{\rm Span}}\{e_{5},e_{6},e_{7}\}.$ By using the Lie bracket in Table 1, it is straightforward to verify that
 $\mathrm{ad}(\mathfrak{so}(3)){\mathfrak{m}}_{1}=0,\quad\mathrm{ad}(\mathfrak{so}(3)){\mathfrak{m}}_{2}\subseteq{\mathfrak{m}}_{2}\quad\text{and}\quad\mathrm{ad}(\mathfrak{so}(3)){\mathfrak{m}}_{3}\subseteq{\mathfrak{m}}_{3}.$
@@
 The subalgebra $\mathfrak{so}(2)\oplus\mathfrak{so}(3)=\operatorname{{\rm Span}}\{e_{1},e_{8},e_{9},e_{10}\}\subset\mathfrak{so}(5)$ provides a reductive decomposition $\mathfrak{so}(5)=\mathfrak{so}(2)\oplus\mathfrak{so}(3)\oplus{\mathfrak{p}}$ (where ${\mathfrak{p}}={\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}$) for the symmetric space
-$G_{2}(\mathbb{R}^{5})=\frac{{\rm SO}(5)}{{\rm SO}(2)\times{\rm SO}(3)},$ (2.5)
+$G_{5,2}=\frac{{\rm SO}(5)}{{\rm SO}(2)\times{\rm SO}(3)}.$ (2.5)
 Furthermore, ${\mathfrak{p}}\simeq\mathbb{C}^{3}$ with respect to the canonical almost complex structure $J=\begin{pmatrix}0&-I_{3}\\ I_{3}&0\end{pmatrix}$ and it turns out that ${\mathfrak{m}}_{3}=\operatorname{{\rm Span}}\{Je_{2},Je_{3},Je_{4}\}$.
 With respect to the basis (2.2), we define the alternating forms on $\mathfrak{p}$} given by
-$\begin{array}[]{rl}\omega_{0}&=e^{25}+e^{36}+e^{47},\\[3.0pt] \mathop{\mathrm{Im}}\Omega_{0}&=e^{237}-e^{246}+e^{345}-e^{567},\\[3.0pt] \mathop{\mathrm{Re}}\Omega_{0}&=e^{234}-e^{267}+e^{357}-e^{456}.\end{array}$ (2.6)
+$\begin{array}[]{rl}\omega_{0}&=e^{25}+e^{36}+e^{47},\\[4.0pt] \mathop{\mathrm{Im}}\Omega_{0}&=e^{237}-e^{246}+e^{345}-e^{567},\\[4.0pt] \mathop{\mathrm{Re}}\Omega_{0}&=e^{234}-e^{267}+e^{357}-e^{456}.\end{array}$ (2.6)
 Notice that $\omega_{0}\wedge\mathop{\mathrm{Re}}\Omega_{0}=0$ and $2\omega_{0}^{3}=3\mathop{\mathrm{Re}}\Omega_{0}\wedge\mathop{\mathrm{Im}}\Omega_{0}\neq 0$, i.e.
-$(\omega_{0},\mathop{\mathrm{Re}}\Omega_{0})$ is a ${\rm SU}(3)$-structure on ${\mathfrak{p}}$ (with induced orientation $-e^{234567}$).
+$(\omega_{0},\mathop{\mathrm{Re}}\Omega_{0})$ is a ${\rm SU}(3)$-structure on $\mathfrak{p}$} (with induced orientation $e^{234567}$).
 Lemma 2.1.
-The forms in (2.6) are $\mathrm{ad}(\mathfrak{so}(2)\oplus\mathfrak{so}(3))$-invariant, therefore, it induces a ${\rm SO}(5)$-invariant ${\rm SU}(3)$-structure $(\omega,\mathop{\mathrm{Re}}\Omega)$ on $G_{2}(\mathbb{R}^{5})$.
+The forms in (2.6) are $\mathrm{ad}(\mathfrak{so}(2)\oplus\mathfrak{so}(3))$-invariant, therefore, they induce a ${\rm SO}(5)$-invariant ${\rm SU}(3)$-structure $(\omega,\mathop{\mathrm{Re}}\Omega)$ on $G_{2}(\mathbb{R}^{5})$.
 Furthermore, we have that $d\omega=0,$ $d(\mathop{\mathrm{Re}}\Omega)=3e^{1}\wedge\mathop{\mathrm{Im}}\Omega$ and $d(\mathop{\mathrm{Im}}\Omega)=-3e^{1}\wedge\mathop{\mathrm{Re}}\Omega.$
@@
 $\mathfrak{stab}_{\mathfrak{gl}({\mathfrak{p}})}(\mathop{\mathrm{Re}}\Omega_{0})=\mathfrak{stab}_{\mathfrak{gl}({\mathfrak{p}})}(\mathop{\mathrm{Im}}\Omega_{0})=\mathfrak{sp}(\mathbb{R}^{6}).$
-Notice that the endomorphisms $\mathrm{ad}(e_{8})_{\mathfrak{p}}$, $\mathrm{ad}(e_{9})_{\mathfrak{p}}$, $\mathrm{ad}(e_{10})_{\mathfrak{p}}$ and $\mathrm{ad}(e_{1})_{\mathfrak{p}}=J$ belong to the interception $\mathfrak{sl}(\mathbb{C}^{3})\cap\mathfrak{sp}(\mathbb{R}^{6})$.
+Notice that the endomorphisms $\mathrm{ad}(e_{8})_{\mathfrak{p}}$, $\mathrm{ad}(e_{9})_{\mathfrak{p}}$, $\mathrm{ad}(e_{10})_{\mathfrak{p}}$ and $\mathrm{ad}(e_{1})_{\mathfrak{p}}=J$ belong to the intersection $\mathfrak{sl}(\mathbb{C}^{3})\cap\mathfrak{sp}(\mathbb{R}^{6})$.
 Therefore, the forms in (2.6) are $\mathrm{ad}(\mathfrak{so}(2)\oplus\mathfrak{so}(3))$-invariant.
@@
 $\omega_{p}={\rm L}_{x}^{\ast}\omega_{0}\quad\text{and}\quad\real\Omega_{p}={\rm L}_{x}^{\ast}\real\Omega_{0},$
-where ${\rm L}_{x}$ is the left multiplication by $x\in{\rm SO}(5)$, hence $2\omega^{3}=3\mathop{\mathrm{Re}}\Omega\wedge\mathop{\mathrm{Im}}\Omega\neq 0$ Finally, from Table 1 follows that $d\omega_{0}=0,$ $d(\mathop{\mathrm{Re}}\Omega_{0})=3e^{1}\wedge\mathop{\mathrm{Im}}\Omega_{0}$ and $d(\mathop{\mathrm{Im}}\Omega_{0})=-3e^{1}\wedge\mathop{\mathrm{Re}}\Omega_{0}$.
+where ${\rm L}_{x}$ is the left multiplication by $x\in{\rm SO}(5)$, hence $2\omega^{3}=3\mathop{\mathrm{Re}}\Omega\wedge\mathop{\mathrm{Im}}\Omega\neq 0$ .
+Finally, from Table 1 follows that $d\omega_{0}=0,$ $d(\mathop{\mathrm{Re}}\Omega_{0})=3e^{1}\wedge\mathop{\mathrm{Im}}\Omega_{0}$ and $d(\mathop{\mathrm{Im}}\Omega_{0})=-3e^{1}\wedge\mathop{\mathrm{Re}}\Omega_{0}$.
 ∎
-The following result is mentioned in [kerr1998new]*§4, for sake of completeness, we include the detail proof of it.
+The following result is mentioned in [kerr1998new]*§4. For the sake of completeness, we include the detailed proof.
 Lemma 2.2.
@@
 $g=\begin{pmatrix}y_{1}&0&0\\ 0&y_{2}I_{3}&xI_{3}\\ 0&xI_{3}&y_{3}I_{3}\end{pmatrix}\quad\text{for}\quad y_{1},y_{2},y_{3}>0,\quad x\in\mathbb{R}\quad\text{with}\quad y_{2}y_{3}>x^{2}.$
-Denote by $N({\rm SO}(3))$ the normalizer of ${\rm SO}(3)$ in ${\rm SO}(5)$, since the quotient $\frac{N({\rm SO}(3))}{{\rm SO}(3)}$ is isomorphic to ${\rm SO}(2)$ the conjugation induces a 1-parameter family of diffeomorphism on ${\mathfrak{m}}$ given by $\phi(t)=\mathrm{Ad}(\exp(te_{1}))$.
+Denote by $N({\rm SO}(3))$ the normalizer of ${\rm SO}(3)$ in ${\rm SO}(5)$, since the quotient $\frac{N({\rm SO}(3))}{{\rm SO}(3)}$ is isomorphic to ${\rm SO}(2)$ the conjugation induces a 1-parameter family of diffeomorphisms on $\mathfrak{m}$} given by $\phi(t)=\mathrm{Ad}(\exp(te_{1}))$.
 Thus, for $t\in\mathbb{R}$ the one-parameter family $g_{t}=\phi(t)^{T}g\phi(t)$ defines an isometric $\mathrm{Ad}({\rm SO}(3))$-inner product.
-A straightforward computation shows that the parameter $x(t)$ in the metric $g_{t}$, is given by $x(t)=\frac{(y_{3}-y_{2})}{2}\sin(2t)+x\cos(2t)$. Hence, $g_{t}$ is diagonal if:
+A direct computation shows that the parameter $x(t)$ in the metric $g_{t}$, is given by $x(t)=\frac{(y_{3}-y_{2})}{2}\sin(2t)+x\cos(2t)$. Hence, $g_{t}$ is diagonal if:
 $y_{2}\neq y_{3}\quad\text{and}\quad t=\frac{1}{2}\arctan\left(\frac{2x}{y_{2}-y_{3}}\right)$
@@
 Lemma 2.3.
-The $5$-dimensional family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$ are described, up to $\mathrm{G}_{2}$-equivalence, by $(a,b,x,y,z)\in\mathbb{R}^{5}$ satisfying
+The $5$-dimensional family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$ is described, up to $\mathrm{G}_{2}$-equivalence, by $(a,b,x,y,z)\in\mathbb{R}^{5}$ satisfying
 $a^{2}+by>0,\quad b^{2}+ax>0\quad z\neq 0\quad\text{and}\quad xy=ab.$ (2.10)
-Explicitly, the $\mathrm{G}_{2}$-structure $\varphi$ can be written by
+Explicitly, the $\mathrm{G}_{2}$-structure $\varphi$ can be written as
 $\varphi=$ $-X^{1}\wedge(X^{25}+X^{36}+X^{47})+p_{1}X^{234}+p_{2}(-X^{267}+X^{357}-X^{456})$ (2.11)
@@
 Proof.
-From (2.6) is easy to check that any $\mathrm{Ad}({\rm SO}(3))$-invariant $3$-form on ${\mathfrak{m}}$ is given by
+From (2.6) it is straightforward to check that any $\mathrm{Ad}({\rm SO}(3))$-invariant $3$-form on $\mathfrak{m}$} is given by
 $\widetilde{\varphi}$= $-\widetilde{z}e^{1}\wedge(e^{25}+e^{36}+e^{47})+\widetilde{x}e^{234}+\widetilde{a}(-e^{267}+e^{357}-e^{456})$
 $+\widetilde{b}(e^{237}-e^{246}+e^{345})-\widetilde{y}e^{567}.$
-The $3$-form $\widetilde{\varphi}$ defines a $\mathrm{G}_{2}$-structure, if the induced symmetric bilinear form
-$B_{ij}=\Big{(}(e_{i}\lrcorner\varphi)\wedge(e_{j}\lrcorner\varphi)\wedge\varphi\Big{)}(e_{1},...,e_{7})$
-is positive (or negative) defined. Explicitly, we obtain that for $k=2,3,4$
-$B_{11}=6\widetilde{z}^{3},\quad B_{kk}=6\widetilde{z}(\widetilde{b}^{2}+\widetilde{a}\widetilde{x}),\quad B_{k+3k+3}=6\widetilde{z}(\widetilde{a}^{2}+\widetilde{b}\widetilde{x}),\quad B_{kk+3}=3\widetilde{z}(\widetilde{a}\widetilde{b}-\widetilde{x}\widetilde{y})$
+The $3$-form $\widetilde{\varphi}$ defines a $\mathrm{G}_{2}$-structure if the induced symmetric bilinear form
+$B_{ij}=\Big((e_{i}\lrcorner\widetilde{\varphi})\wedge(e_{j}\lrcorner\widetilde{\varphi})\wedge\widetilde{\varphi}\Big)(e_{1},...,e_{7})$
+is positive (or negative) definite. Explicitly, we obtain that $B_{11}=6\widetilde{z}^{3}$ and for $k=2,3,4$
+[…]
 and $B_{ij}=0$ otherwise.
@@
 satisfies (2.10). On the other hand, the induced metric is given by
-$g=\frac{B}{6^{\frac{2}{9}}\mathop{\mathrm{missing}}{det}\nolimits(B)^{\frac{1}{9}}},$
-hence, by comparing $g$ with (2.7), we obtain the condition in (2.13). Finally, the expression in (2.11) is obtained by writing $\varphi$ in the orthonormal basis (2.8). ∎
+$g=\frac{B}{6^{\frac{2}{9}}\mathop{\mathrm{det}}\nolimits(B)^{\frac{1}{9}}},$
+hence, by comparing $g$ with (2.7), we obtain the conditions in (2.13). Finally, the expression in (2.11) is obtained by writing $\varphi$ in the orthonormal basis (2.8). ∎
 Definition 2.4.
@@
 Among the class of coclosed class, $\varphi$ is said purely coclosed if $d\varphi\wedge\varphi=0$ and nearly parallel if $d\varphi=4\psi$.
-From Lemma (2.3), the following result describe all coclosed invariant $\mathrm{G}_{2}$–metrics on $V^{5,2}$
+From Lemma (2.3), the following result describes all coclosed invariant $\mathrm{G}_{2}$–metrics on $V^{5,2}$
 Theorem 2.1.
@@
 $+\frac{3p_{2}y_{2}}{z}X^{1567}+\frac{3p_{3}y_{3}}{z}X^{1234}-\frac{p_{4}y_{3}+2p_{3}y_{2}}{z}\left(X^{1267}-X^{1357}+X^{1456}\right).$
-From a straightforward computation, we note that $\ast\left(d\varphi\wedge\varphi\right)=12+\frac{3}{y_{2}}+\frac{3}{y_{3}}\neq 0,$ in other words, $\varphi$ given in (2.11) is not purely coclosed.
+From a direct computation, we note that $\ast\left(d\varphi\wedge\varphi\right)=12+\frac{3}{y_{2}}+\frac{3}{y_{3}}\neq 0,$ in other words, $\varphi$ given in (2.11) is not purely coclosed.
 Finally, by imposing the condition $d\varphi=4\psi$, and using (2.15) and (2.17), we obtain (2.14).
@@
 Remark 2.5.
-The $S^{1}$-family given in (2.14) correspond to the general construction given [alexandrov2012]*(4.28).
+The $S^{1}$-family given in (2.14) corresponds to the general construction given in [alexandrov2012]*(4.28).
 2.2 Homogeneous Sasakian metrics on $V^{5,2}$
 In this section, we study the Sasakian condition for ${\rm SO}(5)$-invariant metrics on $V^{5,2}$.
-Firstly, we characterise the class $(\eta,\xi,g,\Phi)$ of contact metrics structures, cf.
+Firstly, we characterise the class $(\eta,\xi,g,\Phi)$ of contact metric structures, cf.
 Section 1.
@@
 First, note that $(V^{5,2},\eta)$ is a contact manifold, since it is clear that $\eta(\xi)=1$ and that $\eta\wedge d\eta^{3}\neq 0$ follows from (2.16).
-Now, we define an $\mathrm{Ad}({\rm SO}(3))$-invariant contact metric structure $(\eta,\xi,g,\Phi)$, i.e., we define an $\mathrm{Ad}({\rm SO}(3)$-invariant endomorphism $\Phi$ on ${\mathfrak{m}}$ such that $\Phi^{2}=-I+\eta\otimes\xi$ and $d\eta(X,Y)=2g(X,\Phi(Y))$.
+Now, we define an $\mathrm{Ad}({\rm SO}(3))$-invariant contact metric structure $(\eta,\xi,g,\Phi)$, i.e., we define an $\mathrm{Ad}({\rm SO}(3))$-invariant endomorphism $\Phi$ on $\mathfrak{m}$} such that $\Phi^{2}=-I+\eta\otimes\xi$ and $d\eta(X,Y)=2g(X,\Phi(Y))$.
 From (2.9) we have
@@
 $\Phi X_{i}=-\frac{1}{2}\sqrt{\frac{y_{1}}{y_{2}y_{3}}}X_{i+3},\quad\Phi X_{i+3}=\frac{1}{2}\sqrt{\frac{y_{1}}{y_{2}y_{3}}}X_{i},\quad\text{and}\quad\Phi\xi=0.$ (2.19)
-Moreover, using the above equalities and the condition $\Phi^{2}=-I+\eta\otimes\xi$, we obtain that $-\frac{1}{4}\frac{y_{1}}{y_{2}y_{3}}=-1$, or equivalently $y_{1}=4y_{2}y_{3}$. ∎
+Moreover, using the above equalities and the condition $\Phi^{2}=-I+\eta\otimes\xi$, we obtain that $\frac{1}{4}\frac{y_{1}}{y_{2}y_{3}}=-1$, or equivalently $y_{1}=4y_{2}y_{3}$. ∎
 The Nijenhuis tensor induced by $\Phi$ is defined by
@@
 $=\frac{1}{\sqrt{y_{1}}}X_{i}-\frac{1}{\sqrt{y_{1}}}X_{i}=0.$
-For the missing cases, for $X,Y\in{\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}$ the brackets $[X,Y]$, $[\Phi Y,Y]$ and $[X,\Phi Y]$ are generated by $\xi$ (cf.
+For the missing cases, for $X,Y\in{\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}$ the brackets $[X,Y]$, $\Phi Y,Y]$ and $X,\Phi Y]$ are generated by $\xi$ (cf.
 (2.9)), hence $\Phi[X,Y]=\Phi[\Phi X,Y]=\Phi[X,\Phi Y]=0$ and thus $N_{\Phi}(X,Y)=[\Phi X,\Phi Y]$.
@@
 $x=a,\quad y=b\quad\text{and}\quad z^{2}=4(a^{2}+b^{2})^{\frac{4}{3}}.$ (2.23)
-In this case, the $2$-parameter family of coclosed $\mathrm{G}_{2}$-structures is
+In this case, the 22-parameter family of coclosed $\mathrm{G}_{2}$-structures is
 $\varphi=$ $-X^{1}\wedge(X^{25}+X^{36}+X^{47})+\left(\frac{2}{z}\right)^{\frac{3}{4}}\Big\{a\Big(X^{234}-X^{267}+X^{357}-X^{456}\Big)$
@@
 Proof.
-To proof item $(1)$, we compare (2.13) and (2.22), thus
+We compare (2.13) and (2.22), thus
 $z^{2}=4(a^{2}+by)^{4/3}\quad\text{and}\quad a^{2}+by=b^{2}+ax>0,$
@@
 $\hat{F}_{A}\colon=\langle F_{A},\omega\rangle=0,\quad\text{and}\quad F^{0,2}_{A}=0.$ (3.1)
-Here, the type $(p,q)$ of a $k$-form on $M$ is induced by the condition $\Phi|_{H}^{2}=-1$, where $H$ is the Kähler distribution $H=\ker\eta\subset TM=H\oplus\xi$.
+Here, the type $(p,q)$ of a $k$-form on $M$ is induced by the condition $\Phi|_{H}^{2}=-1$, where $H$ is the Kähler distribution given by $H=\ker\eta\subset TM=H\oplus\xi$.
 Additionally, the splitting on the tangent bundle induced by the contact structure yields the splitting $\Omega^{k}(M)=\Omega^{k}(H^{\ast})\oplus\eta\wedge\Omega^{k-1}(H^{\ast})$.
 Hence, after complexification of the tangent bundle, one can naturally extend this bi-degree $(p,q)$ to the bundle of $E$-valued $k$-forms.
-We know that $2$-forms on $M$, and consequently $\Omega^{2}({\mathfrak{g}}_{E})$, decomposes by (cf.
-[portilla2023, Section.
-2.1])
+We know that $2$-forms on $M$, and consequently $\Omega^{2}({\mathfrak{g}}_{E})$, decomposes as (cf. [portilla2023, Section. 2.1])
 $\Omega^{2}(M)=\underbrace{\Omega^{2}_{1}\oplus\Omega^{2}_{6}\oplus\Omega^{2}_{8}}_{\text{Transversal $2$-forms}}\oplus\Omega^{2}_{V}$ (3.2)
@@
 Lemma 3.1.
-With respect to the Sasakian structure of $V^{5,2}$ given in Theorem 2.2 and the $\mathrm{G}_{2}$-structure in (2.11) we have that: If $A$ is a SDCI, then it is a $\mathrm{G}_{2}$-instanton.
+With respect to the Sasakian structure of $V^{5,2}$ given in Theorem 2.2 and the $\mathrm{G}_{2}$-structure in Corollary 2.8 we have that: If AA is a SDCI, then it is a $\mathrm{G}_{2}$-instanton.
 Proof.
-We know that in local coordinates, a basis for self dual contact $2$–forms is given by (cf. [portilla2023, Section 2])
+We know that in local coordinates, a basis for self dual contact 22–forms is given by (cf. [portilla2023, Section 2])
 $\begin{array}[]{lccl}w_{1}=X^{23}+X^{56},&w_{2}=X^{26}+X^{35},&w_{3}=X^{24}+X^{57},&w_{4}=X^{27}+X^{45},\\[4.0pt] w_{5}=X^{34}+X^{67},&w_{6}=X^{37}+X^{46},&w_{7}=X^{25}-X^{47},&w_{8}=X^{36}-X^{47}.\end{array}$ (3.3)
@@
 Thus, for the $\mathrm{G}_{2}$-structure (cf.
-Corollary 2.8):
+Corollary 2.8)
 $\varphi=-\eta\wedge d\eta+\left(\frac{2}{z}\right)^{\frac{3}{4}}a\mathop{\mathrm{Re}}(\Omega_{0})+\left(\frac{2}{z}\right)^{\frac{3}{4}}b\mathop{\mathrm{Im}}(\Omega_{0}),$
@@
 where $\Omega^{2}_{8}\simeq\mathfrak{su}(3)$ and $\Omega^{2}_{1}=\Omega^{0}(M)\otimes\omega$.
-On the other hand, we know that the space of $2$-forms on $H^{\mathbb{C}}=H\otimes_{\mathbb{R}}\mathbb{C}$ decomposes into $\operatorname{Gl}(3,\mathbb{C})$-invariant submodules
+On the other hand, we know that the space of $2$-forms on $H^{\mathbb{C}}=H\otimes_{\mathbb{R}}\mathbb{C}$ decomposes into the $\operatorname{Gl}(3,\mathbb{C})$-invariant submodules
 $\Omega^{2}(H^{\mathbb{C}})=\Omega^{2,0}(H^{\mathbb{C}})\oplus\Omega^{0,2}(H^{\mathbb{C}})\oplus\Omega^{1,1}(H^{\mathbb{C}}).$
@@
 Since ${\rm SU}(3)\subset\mathrm{G}_{2}$ we have that $\Omega_{8}^{2}\subset\Omega_{14}^{2}$ and $\Omega^{2}_{1}\subset\Omega^{2}_{7}$.
-It is easy to see that $\Omega^{2}_{V}$ has elements in both $\Omega_{7}^{2}$ and $\Omega_{14}^{2}$, hence the same is true for $\Omega^{2}_{6}$.
+It is easy to see that $\Omega^{2}_{V}$ has elements in both $\Omega_{7}^{2}$ and $\Omega_{14}^{2}$.
+Hence, the same is true for $\Omega^{2}_{6}$.
 Lemma 3.3.
-Let $E\to V^{5,2}$ be a Sasakian holomorphic vector bundle [Boyer2008, Section 3], with the Sasakian structure $(\eta,\xi,g,\Phi)$ given in Theorem 2.2 and $A\in\mathcal{A}(E)$ the Chern connection.
-If $A\in\mathcal{A}(E)$ is a $\mathrm{G}_{2}$-instanton then $A$ satisfies the equation (1.2), i.e.
+Let $E\to V^{5,2}$ be a Sasakian holomorphic vector bundle [Boyer2008, Section 3], the Sasakian structure $(\eta,\xi,g,\Phi)$ is given in Theorem 2.2 and $A\in\mathcal{A}(E)$ the Chern connection.
+If $A\in\mathcal{A}(E)$ is a $\mathrm{G}_{2}$-instanton, then $A$ satisfies the equation (1.2), i.e.
 it is a $\lambda$-contact instanton .
@@
 $=F_{A}\wedge\eta\wedge d\eta.$
-Consequently, solutions to the $\mathrm{G}_{2}$-instanton equation are solutions to the contact instanton equation. ∎
+Consequently, solutions to the $\mathrm{G}_{2}$-instanton equation are also solutions to the contact instanton equation. ∎
 Consider the Kähler manifold $G_{2}(\mathbb{R}^{5})\cong V^{5,2}/S^{1}$, we fix a holomorphic bundle $(\mathcal{E},\bar{\partial})$ on $G_{5,2}$, the pullback $\mathcal{E}^{\ast}:=\pi^{\ast}\mathcal{E}\to V^{5,2}$ carries a structure of Sasakian holomorphic bundle
@@
 The vector bundle $\mathcal{E}^{\ast}$ is a holomorphic Sasakian bundle, and a Chern connection $A$ on $\mathcal{E}$ is HYM if, and only if, $\pi^{\ast}A$ is a $\mathrm{G}_{2}$-instanton on $\mathcal{E}^{*}$.
-We recall that in Lemma 3.1, Lemma 3.3 and Corollary 3.4 the key argument follow just in presence of a SU$(n)$-structure.
+We recall that in Lemma 3.1, Lemma 3.3 and Corollary 3.4 the key argument follow just in presence of a SU(n)(n)-structure.
 These have analogous results in [portilla2023, Proposition 3.11] and [calvo2020gauge, Lemma 21], the difference in our case is that we do not assume $M$ is a contact Calabi-Yau but only endowed with a transversal ${\rm SU}(3)$-structure.
 4 Homogeneous Principal Bundles on $V^{5,2}$
-First, we recall some notation and results about homogeneous principal bundle, following [cap2009]*Section 1.4.
-Let $H\subset G$ be a closed subgroup of the Lie group $G$, we denote by $M=G/H$ the homogeneous space, $L\colon G\times M\to M$ the canonical left action and $p\colon G\to M$ an $H$–principal bundle given by the canonical projection.
+First, we recall some notation and results about homogeneous principal bundles, following [cap2009]*Section 1.4.
+Let $H\subset G$ be a closed subgroup of the Lie group $G$, we denote by $M=G/H$ the homogeneous space, let $L\colon G\times M\to M$ be the canonical left action and $p\colon G\to M$ an HH–principal bundle given by the canonical projection.
 Definition 4.1.
-A homogeneous principal bundle over $M$ is a locally trivial principal bundle $\pi\colon P\to M$ together with a left $G$–action $\widetilde{L}\colon G\times P\to P$, which lifts the action $L$ on $M$, i.e., $\pi(\widetilde{L}({\textsl{g}},y))=L({\textsl{g}},\pi(y))$, and for each ${\textsl{g}}\in G$, the bundle map $\widetilde{L}_{\textsl{g}}\colon P\to P$ is a homomorphism of principal bundles.
+A homogeneous principal bundle over $M$ is a locally trivial principal bundle $\pi\colon P\to M$ together with a left GG–action $\widetilde{L}\colon G\times P\to P$, which lifts the action $L$ on $M$, i.e., $\pi(\widetilde{L}({\textsl{g}},y))=L({\textsl{g}},\pi(y))$, and for each ${\textsl{g}}\in G$, the bundle map $\widetilde{L}_{\textsl{g}}\colon P\to P$ is a homomorphism of principal bundles.
 Homogeneous principal bundles are characterised in the following way:
@@
 [cap2009, Lemma 1.4.5] Let $G$ and $K$ be Lie groups and $H\subset G$ a closed Lie subgroup.
-Let $P\to G/H$ a homogeneous $K$-principal bundle, then there exist a smooth homomorphism $\phi\colon H\to K$ such that $P\cong G\times_{\phi}K$.
+Let $P\to G/H$ be a homogeneous $K$-principal bundle, then there exists a smooth homomorphism $\phi\colon H\to K$ such that $P\cong G\times_{\phi}K$.
 The action of $H$ on $K$ is given by $h\cdot k=\phi(h)\cdot k$ for each $h\in H$ and $k\in K$.
-The bundles corresponding to two homomorphisms $\phi,\widetilde{\phi}\colon H\to K$ are isomorphic if and only if there exist $k\in K$ such that $k\cdot\phi=\widetilde{\phi}\cdot k$
-For a homogeneous $K$-principal bundle $P\to G/H$, a connection $A\in\Omega^{1}(P,{\mathfrak{k}})$ is called invariant, if $\widetilde{L}_{\textsl{g}}^{*}A=A$ for each ${\textsl{g}}\in G$.
+The bundles corresponding to two homomorphisms $\phi,\widetilde{\phi}\colon H\to K$ are isomorphic if and only if there exists $k\in K$ such that $k\cdot\phi=\widetilde{\phi}\cdot k$
+For a homogeneous $K$-principal bundle $P\to G/H$, a connection $A\in\Omega^{1}(P,{\mathfrak{k}})$ is called invariant if $\widetilde{L}_{\textsl{g}}^{*}A=A$ for each ${\textsl{g}}\in G$.
 Lemma 4.3.
@@
 Since ${\rm SO}(3)$ is simple, up to conjugation there are two smooth homomorphisms $\phi_{i}\colon{\rm SO}(3)\to{\rm SO}(3),$ $i=0,1$.
-Namely, $\phi_{0}$ the trivial homomorphism and $\phi_{1}=1_{{\rm SO}(3)}$ the identity, according to Lemma 4.2, there exist precisely two [^3] Up to isomorphism of homogeneous bundles homogeneous ${\rm SO}(3)$-principal bundles $\pi_{i}\colon P_{i}=G\times_{\phi_{i}}{\rm SO}(3)\to V^{5,2}$, $i=0,1$.
+Namely, $\phi_{0}$ the trivial homomorphism and $\phi_{1}=1_{{\rm SO}(3)}$ the identity homomorphism, according to Lemma 4.2, there exist precisely two [^3] Up to isomorphism of homogeneous principal bundles homogeneous ${\rm SO}(3)$-principal bundles $\pi_{i}\colon P_{i}=G\times_{\phi_{i}}{\rm SO}(3)\to V^{5,2}$, $i=0,1$.
 Lemma 4.4.
@@
 Now, we apply the second condition to ${\textsl{g}}_{k}=\begin{pmatrix}I_{2\times 2}&\\ 0&h_{k}\\ \end{pmatrix}\in{\rm SO}(5),$ we define some values of $h_{k}$,
-$h_{0}=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&-1\end{pmatrix},h_{1}=\begin{pmatrix}0&1&0\\ -1&0&0\\ 0&0&1\end{pmatrix},h_{2}=\begin{pmatrix}0&0&1\\ 0&-1&0\\ 1&0&0\end{pmatrix},\quad\text{and}\quad h_{3}=\begin{pmatrix}0&0&1\\ 0&1&0\\ -1&0&0\end{pmatrix}$ (4.2)
+$h_{0}=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&-1\end{pmatrix},h_{1}=\begin{pmatrix}0&1&0\\ -1&0&0\\ 0&0&1\end{pmatrix},h_{2}=\begin{pmatrix}0&0&1\\ 0&-1&0\\ 1&0&0\end{pmatrix}\quad\text{and}\quad h_{3}=\begin{pmatrix}0&0&1\\ 0&1&0\\ -1&0&0\end{pmatrix}$ (4.2)
 First, considering $h_{0}$ and $h_{1}$, we obtain that $\alpha(e_{i})=a_{i}e_{8}+b_{i}e_{9}+c_{i}e_{10}$ and, for $i\neq 3$ and $i\neq 6$
@@
 $\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{i}))=\alpha({\textsl{g}}_{1}e_{i}{\textsl{g}}_{1}^{-1})$ $=-a_{i}e_{8}-b_{i}e_{10}-c_{i}e_{9}$
-follows that $a_{i}=b_{i}=c_{i}=0$ for $i\in\{2,4,5,7\}$. Now, by considering $h_{2}$ and $h_{3},$ we obtain for $i=3,6$
+it follows that $a_{i}=b_{i}=c_{i}=0$ for $i\in\{2,4,5,7\}$. Now, by considering $h_{2}$ and $h_{3},$ we obtain for $i=3,6$
 $\begin{array}[]{cl}\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{i}))=&-\alpha(e_{i})=a_{i}e_{8}+b_{i}e_{10}+c_{i}e_{9}\\[5.0pt] \alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{i}))=&\alpha(e_{i})=-a_{i}e_{8}-b_{i}e_{10}-c_{i}e_{9}.\end{array}$
@@
 $\alpha=e^{1}\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10}).$ (4.3)
-The curvature of $\alpha$ is $F_{\alpha}=y_{2}^{-1}(X^{25}+X^{36}+X^{47})\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10})$. The second affirmation in the Lemma follows from (2.6) since
+The curvature of $\alpha$ is $F_{\alpha}=y_{2}^{-1}(X^{25}+X^{36}+X^{47})\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10})$. The second statement in the Lemma follows from (2.6) since
 $F_{\alpha}=2\omega\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10}).$ (4.4)
-Remark 4.5 (Homogeneous ${\rm U}(1)$-connections).
+Lemma 4.5 (Homogeneous ${\rm U}(1)$-connections).
+There is a unique invariant ${\rm U}(1)$-connections on $P=V^{5,2}\times{\rm U}(1)$, given by
+$\beta=a_{1}e^{1}\otimes e_{1}$ (4.5)
+Furthermore, the curvature $F_{\alpha}$ of the connection $\alpha$ belongs to $\Omega^{2}_{1}$ cf. (3.4).
+Proof.
 Similar to Lemma 4.4, since ${\rm SO}(3)$ is simple, all non-trivial homomorphism $\phi\colon{\rm SO}(3)\rightarrow K$ is either trivial or injective for all Lie group $K$.
@@
 $\alpha=a_{i}e^{i}\otimes e_{1}\quad\text{for}\quad i=1,\dots,10.$
-From item $(i)$, since $\alpha|_{\mathfrak{so}(3)}=0$, $\alpha$ has to be of the form $\alpha=a_{i}e^{i}\otimes e_{1}$, $i=1,\cdots,7.$ On the other hand, considering item $(ii)$ for ${\textsl{g}}_{0},{\textsl{g}}_{1}$ given in (4.2), we use the same argument used in the proof of Lemma 4.4 to obtain $\alpha(e_{i})=a_{i}e_{1}$ and
-$\alpha(Ad({\textsl{g}}_{0})(e_{i}))=\alpha({\textsl{g}}_{0}e_{i}{\textsl{g}}_{0}^{-1}))$ $=-a_{i}e_{1}$
-$\alpha(Ad({\textsl{g}}_{1})(e_{i}))=\alpha({\textsl{g}}_{1}e_{i}{\textsl{g}}_{1}^{-1}))$ $=a_{i}e_{1}.$
-for $i=2,\cdots,7$. Therefore, $\alpha$ is characterised by
+From item $(i)$, since $\alpha|_{\mathfrak{so}(3)}=0$, $\alpha$ has to be of the form $\alpha=a_{i}e^{i}\otimes e_{1}$, $i=1,\cdots,7.$ On the other hand, considering item $(ii)$ for ${\textsl{g}}_{0},{\textsl{g}}_{1}$ given in (4.2), we use the same argument used in the proof of Lemma 4.4 to obtain $\alpha(e_{i})=a_{i}e_{1}$ and for $i=2,\cdots,7$
+$\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{i}))=\alpha({\textsl{g}}_{0}e_{i}{\textsl{g}}_{0}^{-1}))$ $=-a_{i}e_{1}$
+$\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{i}))=\alpha({\textsl{g}}_{1}e_{i}{\textsl{g}}_{1}^{-1}))$ $=a_{i}e_{1}.$
+Therefore, $\alpha$ is characterised by
 $\beta=a_{1}e^{1}\otimes e_{1}$ (4.6)
@@
 $F_{\beta}=2(X^{25}+X^{36}+X^{47})\otimes e_{1}.$ (4.7)
-Note that $F_{\beta}\in\Omega^{2}_{1}=\omega\otimes\Omega^{0}(M)$ (4.13).
+Note that $F_{\beta}\in\Omega^{2}_{1}=\omega\otimes\Omega^{0}(M)$ (4.14). ∎
 Now, we consider the ${\rm SO}(3)$-principal bundle $P_{1}={\rm SO}(5)\times_{\phi_{1}}{\rm SO}(3)$ induced by the homomorphism identity $\phi_{1}\colon{\rm SO}(3)\to{\rm SO}(3)$.
@@
 $\pi_{1}\colon[(u,s)]\in P_{1}\mapsto u\cdot{\rm SO}(3)\in V^{5,2},$ (4.8)
-and the class is given by $[(u,s)]=\{(u\cdot r,r^{-1}s)\;|\;r\in{\rm SO}(3)\}.$ There exist a left ${\rm SO}(5)$-action on $P_{1}$ such that for each $v\in{\rm SO}(5)$ the bundle map $L_{v}\colon P_{1}\to P_{1}$ is $\pi_{1}$-equivariant i.e.
+and the class is given by $(u,s)]=\{(u\cdot r,r^{-1}s)\;|\;r\in{\rm SO}(3)\}.$ There exist a left ${\rm SO}(5)$-action on $P_{1}$ such that for each $v\in{\rm SO}(5)$ the bundle map $L_{v}\colon P_{1}\to P_{1}$ is $\pi_{1}$-equivariant, i.e.,
 $\pi_{1}(L_{v}[(u,s)])=\pi_{1}([(vu,s)])=vu\cdot{\rm SO}(3)=L_{v}(\pi_{1}[(u,s)]).$
-According to (2.2), we fix the basis $\{e_{8},e_{9},e_{10}\}$ for $\mathfrak{so}(3)\subset\mathfrak{so}(5)$ such that $[e_{i},e_{j}]=\varepsilon_{ijk}e_{k}$, where $\varepsilon_{ijk}$ is the sign of the permutation $(i\;j\;k)$.
+According to (2.2), we fix the basis $\{e_{8},e_{9},e_{10}\}$ for $\mathfrak{so}(3)\subset\mathfrak{so}(5)$ such that $e_{i},e_{j}]=\varepsilon_{ijk}e_{k}$, where $\varepsilon_{ijk}$ is the sign of the permutation (ijk)(i\;j\;k).
 Lemma 4.6.
@@
 Consider $\alpha:\mathfrak{so}(5)\rightarrow\mathfrak{so}(3)$ in the basis (2.2) given by $\alpha=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$.
-From Lemma 4.3 (i), we have that $\alpha=\alpha_{0}+e^{8}\otimes e_{8}+e^{9}\otimes e_{9}+e^{10}\otimes e_{10},$ where $\alpha_{0}=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$, $i,j,k\in\{1,\dots,7\}$, we are going to check that $\alpha_{0}=0$.
+From Lemma 4.3 (i), we have that $\alpha=\alpha_{0}+e^{8}\otimes e_{8}+e^{9}\otimes e_{9}+e^{10}\otimes e_{10},$ where $\alpha_{0}=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$, with $i,j,k\in\{1,\dots,7\}$, we are going to conclude that $\alpha_{0}=0$.
 From Lemma 4.3 (ii), for any ${\textsl{g}}\in{\rm SO}(3)$ and $X\in\mathfrak{so}(5)$, we have ${\textsl{g}}\alpha(X){\textsl{g}}^{-1}=\alpha({\textsl{g}}X{\textsl{g}}^{-1})$.
@@
 $\mathrm{Ad}({\textsl{g}}_{0})(\alpha(e_{2}))=-a_{2}e_{8}+b_{2}e_{10}-c_{2}e_{9}.$
-These imply $-a_{2}=a_{3},-c_{2}=b_{3},c_{3}=b_{2}$. Analogously, for $h_{1}$, we have
+These imply $a_{2}=a_{3},-c_{2}=b_{3},c_{3}=b_{2}$. Analogously, for $h_{1}$, we have
 $\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{2}))=\alpha(e_{3})=-a_{3}e_{8}-b_{3}e_{9}-c_{3}e_{10}$
@@
 $\mathrm{Ad}({\textsl{g}}_{1})(\alpha(e_{2}))=a_{2}e_{8}+b_{2}e_{10}-c_{2}e_{9},$
-consequently, $-a_{2}=a_{3},c_{2}=b_{3},-b_{2}=c_{3}$, hence $c_{2}=b_{3}=c_{3}=b_{2}=0$. Similarly:
+consequently, $a_{2}=a_{3},c_{2}=b_{3},-b_{2}=c_{3}$, hence $c_{2}=b_{3}=c_{3}=b_{2}=0$. Similarly:
   • From $\mathrm{Ad}({\textsl{g}}_{0})(\alpha(e_{4}))=\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{4}))$, we obtain $c_{4}=b_{4}=0$,
@@
 Proof.
-For any Sasakian manifold, the bundle of $2$–forms decomposes into [portilla2023, Section 2.1]
+For any Sasakian manifold, the bundle of 22–forms decomposes into [portilla2023, Section 2.1]
 $\Omega^{2}=\Omega^{2}_{1}\oplus\Omega^{2}_{6}\oplus\Omega^{2}_{8}\oplus\Omega^{2}_{V},$ (4.14)
@@
 Consider $V^{5,2}$ endowed with the Sasakian structure given in Corollary 2.8.
-There exist an unique invariant ASDCI on the trivial homogeneous ${\rm SO}(3)$–principal bundle $P_{0}\to V^{5,2}$ (cf.
+There exist a unique invariant ASDCI on the trivial homogeneous ${\rm SO}(3)$–principal bundle $P_{0}\to V^{5,2}$ (cf.
 Lemma 4.4).
@@
 Proposition 4.9.
-Consider the connection (4.8) on $V^{5,2}$, with respect to the $\mathrm{G}_{2}$–structure given in (2.11). Then we gave
+Consider the connection (4.9) on $V^{5,2}$, with respect to the $\mathrm{G}_{2}$–structure given in Corollary (2.8). Then, we have
   • The ${\rm SO}(3)$–connection in (4.9) on the homogeneous principal bundle $P_{1}\to V^{5,2}$ [cf.
-Lemma 4.6] satisfies $F_{\alpha}\wedge\psi=0$ for any $y_{1}>0$, i.e., there exist an unique ${\rm SO}(5)$–invariant connection which is simultaneously $\mathrm{G}_{2}$ and SDCI, given by (4.11).
+Lemma 4.6] satisfies $F_{\alpha}\wedge\psi=0$ for any $y_{1}>0$, i.e., there exist a unique ${\rm SO}(5)$–invariant connection which is simultaneously $\mathrm{G}_{2}$ and SDCI, given by (4.12).
   • The ${\rm SO}(3)$–connection (4.1) on the trivial bundle $P_{0}\to V^{5,2}$ (cf.
 Lemma 4.4), whose curvature is given in (4.4) satisfies $F_{\alpha}\wedge\psi=0$, for $\psi$ given in (2.15), if and only if (4.1) is flat.
-4.3 Stability of SDCI and $\mathrm{G}_{2}$-instantons
-Now, we analyse the Yang-Mills condition for instanton (4.8), for a detailed study on stability of Yang-Mills connections we refer to [bourguignon1981stability].
-On the space of connections $\mathcal{A}$, the global inner product $(\alpha,\beta)=\int_{M}\langle\alpha\wedge\ast\beta\rangle$ allow us define the Yang-Mills functional as the $L^{2}$-norm of the curvature:
+4.3 Yang–Mills connections
+Now, we analyse the Yang–Mills condition for instanton (4.9), for a detailed study on stability of Yang–Mills connections we refer to [bourguignon1981stability].
+On the space of connections $\mathcal{A}$, the global inner product $(\alpha,\beta)=\int_{M}\langle\alpha\wedge\ast\beta\rangle$ allow us to define the Yang–Mills functional as the $L^{2}$-norm of the curvature:
 $\mathcal{YM}\colon\nabla\in\mathcal{A}\to\|{\rm F}_{\nabla}\|^{2}_{L^{2}}=\int_{M}\operatorname{Tr}({\rm F}_{\nabla}\wedge\ast{\rm F}_{\nabla})=\int_{M}\|{\rm F}_{\nabla}\|^{2}\in\mathbb{R}.$ (4.15)
-A critical point of (4.14) is called a Yang-Mills (YM) connection, i.e.
-$\nabla$ satisfies the so-called Yang-Mills equation $d^{\ast}_{\nabla}F_{\nabla}=0$.
-Taking the exterior derivative $d_{\nabla}$ in (1.2) and using the Bianchi identity, one note that a solution of the instanton equation (1.2) satisfy the YM-equation with torsion:
+A critical point of (4.15) is called a Yang–Mills (YM) connection, i.e.
+$\nabla$ satisfies the so-called Yang–Mills equation $d^{\ast}_{\nabla}F_{\nabla}=0$.
+Taking the exterior derivative $d_{\nabla}$ in (1.2) and using the Bianchi identity, one notes that a solution of the instanton equation (1.2) satisfy the YM-equation with torsion:
 $d_{\nabla}(\ast F_{\nabla})=d\sigma\wedge F_{\nabla}.$
-Note that neither $\sigma=\eta\wedge d\eta$ nor $\sigma=\varphi$ are closed, consequently, the torsion term in the above equation is no automatically zero, however, we have that SDCI and ASDCI are Yang-Mills connections [luis2024Weitzenbock, Section 5].
+Note that neither $\sigma=\eta\wedge d\eta$ nor $\sigma=\varphi$ are closed, consequently, the torsion term in the above equation is not automatically zero, however, we have that SDCI and ASDCI are Yang–Mills connections [luis2024Weitzenbock, Section 5].
 With respect to the $\mathrm{G}_{2}$-structure (2.23), we have the following:
@@
 Let $\alpha$ be the unique $\mathrm{G}_{2}$-instanton (4.9) on the homogeneous ${\rm SO}(3)$-principal bundle $P_{1}$.
-Then, $\alpha$ is a critical point for the Yang-Mills energy for any coclosed $\mathrm{G}_{2}$-structure (2.23).
+Then, $\alpha$ is a critical point for the Yang–Mills energy for any coclosed $\mathrm{G}_{2}$-structure (2.23).
 Proof.
@@
 we see that both equations hold for (2.23), i.e. $x=a$ and $y=b$. Hence, we obtain the result. ∎
-For the Sasakian case, we can prove that SDCI are local minimizers of the YM-functional (4.14).
-In order to prove this, note that the contact structure induces an orthogonal decomposition (cf.
-[portilla2023, Lemma 2.3]) of the curvature $2$-form ${\rm F}_{\nabla}$ into
-${\rm F}_{H}=i_{\xi}(\eta\wedge{\rm F}_{\nabla})\quad\text{and}\quad{\rm F}_{V}=\eta\wedge i_{\xi}{\rm F}_{\nabla},$
-where $F_{H}$ is the horizontal part and $F_{V}$ the vertical part of $F_{\nabla}$. Consequently, we have $\|{\rm F}_{\nabla}\|^{2}=\|{\rm F}_{H}\|^{2}+\|{\rm F}_{V}\|^{2}$ and denote
-${\rm F}_{H}^{+}=\frac{1}{2}({\rm F}_{H}+\ast(\sigma\wedge{\rm F}_{H}))\quad\text{and}\quad{\rm F}_{H}^{-}=\frac{1}{2}({\rm F}_{H}-\ast(\sigma\wedge{\rm F}_{H})),$ (4.15)
-According to the splitting (3.2), we have ${\rm F}_{H}={\rm F}_{8}+{\rm F}_{6}+{\rm F}_{1}\in\Omega_{8}^{2}({\mathfrak{g}}_{E})\oplus\Omega_{6}^{2}({\mathfrak{g}}_{E})\oplus\Omega_{1}^{2}({\mathfrak{g}}_{E})$, then we observe that ${\rm F}_{H}^{+}={\rm F}_{8}-\frac{1}{2}{\rm F}_{1}$ and ${\rm F}_{H}^{-}={\rm F}_{6}+\frac{3}{2}{\rm F}_{1}$, hence
-$\|{\rm F}_{H}\|^{2}$ $=\langle{\rm F}_{H}^{+}+{\rm F}_{H}^{-},{\rm F}_{H}^{+}+{\rm F}_{H}^{-}\rangle=\|{\rm F}_{H}^{+}\|^{2}+\|{\rm F}_{H}^{-}\|^{2}+2\langle{\rm F}_{H}^{+},{\rm F}_{H}^{-}\rangle$
-$=\|{\rm F}_{H}^{+}\|^{2}+\|{\rm F}_{H}^{-}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2},$
-consequently, the YM-functional (4.14) splits as
-$\mathcal{YM}(\nabla)=\|{\rm F}_{H}^{+}\|^{2}+\|{\rm F}_{H}^{-}\|^{2}+\|{\rm F}_{V}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2}.$ (4.16)
-In the Sasakian context, the charge of the connection $\nabla$ is defined by
-$\kappa(\nabla)=\displaystyle\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}^{2})\wedge\eta\wedge\omega.$ (4.17)
-Notice that the charge (4.17) is determined by the horizontal part of $F_{\nabla}$ [portilla2023, Proposition 3.13]:
-$\kappa(\nabla)=\|{\rm F}_{H}^{+}\|^{2}-\|{\rm F}_{H}^{-}\|^{2}.$ (4.18)
-Using (4.16) and (4.18), we have that the YM-mills action is given by
-$\mathcal{YM}(\nabla)$ $=\kappa(\nabla)+2\|{\rm F}_{H}^{-}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2}+\|{\rm F}_{V}\|^{2}$
-$=-\kappa(\nabla)+2\|{\rm F}_{H}^{+}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2}+\|{\rm F}_{V}\|^{2},$
-hence, we have
-$\mathcal{YM}(\nabla)=\begin{cases}\kappa(\nabla),&\nabla\in\Omega^{2}_{8}({\mathfrak{g}}_{E})\\ -\kappa(\nabla),&\nabla\in\Omega^{2}_{6}({\mathfrak{g}}_{E})\\ 2\kappa(\nabla),&\nabla\in\Omega^{2}_{1}({\mathfrak{g}}_{E}).\end{cases}$ (4.19)
-We recall that if $M$ is a contact manifold, a $k$-form $\alpha\in\Omega^{k}(M)$ is called transversal form if $i_{\xi}\alpha=0$, we denote the space of transversal forms by $\Omega^{k}_{H}(M)$.
-Consequently, the space of $k$-forms splits as
-$\Omega^{k}(M)=\Omega^{k}_{H}(M)\oplus\eta\wedge\Omega^{k-1}_{H}(M).$ (4.20)
-More generally, consider a Riemannian manifold $M$ endowed with a Riemannian foliation $\mathcal{F}\subset M$ and consider a principal $G$-bundle $\pi\colon P\to M$.
-$P$ is said to be a foliated principal bundle if there exists a $G$-invariant foliation $\widetilde{\mathcal{F}}$ on $P$ such that the tangent distribution $T\widetilde{\mathcal{F}}\subset TP$ projects isomorphically onto $T\mathcal{F}$ under $\pi_{\ast}\colon TP\to TM$.
-Accordingly, a vector bundle is called foliated if its associated principal bundle is foliated.
-A local section $s\colon U\subset M\to P$ is said basic section if $s_{\ast}\colon TU\to TP$ maps $T\mathcal{F}$ to $T\widetilde{\mathcal{F}}$.
-It is known that every foliated principal bundle admits local basic sections [molino1988riemannian, Proposition 2.7].
-We denote the space of basic sections by $\Gamma(E)_{B}\subset\Gamma(E)$.
-In the context of contact manifolds, $\mathcal{F}$ is naturally given by the Reeb foliation, if $\nabla\in\Gamma(E)$ is a connection with $i_{\xi}F_{\nabla}=0$, then the horizontal lift of $\mathcal{F}$ with respect to $\nabla$ is an integrable distribution and gives $E$ a foliated structure for which $\nabla$ is basic connection [molino1988riemannian], consequently, we have a natural split of the bundle of $E$-valued $k$-forms analogous to (4.20).
-On the underlying bundle $E$, considering only transverse deformations, i.e., $\nabla^{\prime}=\nabla+\alpha$, where $\alpha\in\Omega^{1}_{H}({\mathfrak{g}}_{E})$, in [portilla2023, Lemma 3.14] is shown that the charge vary by
-$\kappa(\nabla)=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\alpha)\wedge\omega^{2}.$ (4.21)
-The following result emulates [portilla2023, Lemma 3.14], the proof proceeds with the same reasoning, but we emphasise that, first, this result is not exclusive to the Stiefel manifold, but rather applies to any Sasakian manifold.
-And secondly, in [portilla2023] the result is announced for transverse deformations of the Chern connection [portilla2023, Corollary 3.15], but actually this result holds in a more general context as follow:
+5 Infinitesimal deformations and rigidity of $\mathrm{G}_{2}$-instantons
+In this section, we extend the spinorial characterisation of infinitesimal deformations of $\mathrm{G}_{2}$–instantons obtained in [singhal2022] to the setting of coclosed $\mathrm{G}_{2}$–structures.
+As an application, we show that for certain values of the parameter $y_{2}$ appearing in (4.12), the corresponding $\mathrm{G}_{2}$–instanton is rigid.
+Throughout this section, $(M^{7},\varphi)$ denotes a 77–dimensional manifold endowed with a coclosed $\mathrm{G}_{2}$–structure, and $\zeta$ the unit spinor associated with $\varphi$.
+Let us recall, consider $\zeta$ the unit spinor associated to $\varphi$ and the induced spinor representation $\Delta_{7}\simeq\langle\zeta\rangle\oplus TM$, the Clifford product of a vector field $Y$ and a spinor $(f,Z)$ is (see [karigiannis2006some])
+$Y\cdot(f,Z)=(-\langle Y,Z\rangle,fY+Y\times Z),$ (5.1)
+where $(f,0)=f\zeta$ and $Y\times Z=(\varphi(Y,Z,\cdot))^{\sharp}$. The Clifford product (5.1) is extended to a $k$-form $\varepsilon$ by
+$\varepsilon\cdot(f,Y)=\frac{1}{k!}\varepsilon_{i_{1},\dots i_{k}}e_{i_{1}}\cdot(e_{i_{2}}\cdot(\dots\cdot(e_{i_{k}}\cdot(f,Z)\dots))$ (5.2)
+Hence, we have the characterisation of the instanton condition (1.3) in the $\mathrm{G}_{2}$ case in terms of spinors:
+Definition 5.1.
+Let $P\to M$ be a principal bundle.
+A connection $A$ on a principal bundle $P$ is called a $\mathrm{G}_{2}$–instanton if its curvature satisfies $F_{A}\cdot\zeta=0,$ where $\cdot$ denotes Clifford multiplication (5.2).
+5.1 Infinitesimal deformation of $\mathrm{G}_{2}$-instantons
+Definition 5.2.
+Let $A$ be a $\mathrm{G}_{2}$-instanton on a principal bundle $P\to M$. An infinitesimal deformation of $A$ is a one–form $\varepsilon\in\Omega^{1}(M,\mathrm{Ad}(P))$ satisfying
+$(d_{A}\varepsilon)\cdot\zeta=0,\qquad d_{A}^{\ast}\varepsilon=0.$ (5.3)
+Let $\nabla^{g}$ be the Levi–Civita connection of $g$ associated to $\varphi$, we consider the one-parameter family of connections
+$\nabla^{t}=\nabla^{g}+\frac{t}{2}\,g^{-1}H_{\varphi}\quad\text{for}\quad t\in\mathbb{R},$ (5.4)
+and
+$H_{\varphi}=\frac{\tau_{0}}{6}\,\varphi-\tau_{3}.$
+For $t=1$, it turns out that $\nabla^{1}$ is the unique connection with parallel skew-symmetric torsion $H_{\varphi}$ that makes $\varphi$ parallel [friedrich2002].
+The connection $\nabla^{t}$ induces a connection on the spinor bundle $\not${S} such that:
+[…]
+Definition 5.3.
+With respect to the Clifford multiplication $\cdot\colon TM\times\not$$\to\not{S}$, we define the Dirac operator
+$D^{t}\mu=\sum_{i=1}^{7}e_{i}\cdot\nabla^{t}_{e_{i}}\mu,$ (5.6)
+[…]
+[…]
+[…]
 Lemma 5.4.
-Let $E$ be a Sasakian ${\rm G}$-bundle over a compact closed 7-manifold.
-For any variation $\nabla^{\prime}=\nabla+\alpha\in\mathcal{A}(E)$, $\alpha\in\Omega^{1}({\mathfrak{g}}_{E})$, on the underlying complex vector bundle E, we have
-$\kappa(\nabla)=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\ast(\omega\otimes s)),$ (4.22)
-where $s\in\Gamma(E)_{B}$ is some transversal section of $E$.
+Let $\zeta$ be the spinor associated with $\varphi$, then it holds
+$(X\lrcorner\varphi)\cdot\zeta=3X\cdot\zeta,\qquad(X\lrcorner\gamma)\cdot\zeta=2h(X)\cdot\zeta\quad\text{and}\quad(X\lrcorner\ast\gamma)\cdot\zeta=-2h(X)\cdot\zeta,$
+where $X\in\mathscr{X}(M)$, $\gamma\in\Omega^{3}_{27}$ and $h$ is the traceless symmetric tensor defining $\gamma$ by
+$\gamma_{ijk}=h_{ip}\gamma_{pjk}+h_{jp}\gamma_{ipk}+h_{kp}\gamma_{ijp}.$
 Proof.
-The proof of [portilla2023, Lemma 3.14] provide that for an arbitrary deformation $\nabla^{\prime}=\nabla+\alpha$, where $\alpha\in\Omega^{1}({\mathfrak{g}}_{E})=\Omega^{1}_{H}({\mathfrak{g}}_{E})\oplus\eta\wedge\Omega^{0}_{H}({\mathfrak{g}}_{E})$, the charge varies as
-$\kappa(\nabla)$ $=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\alpha)\wedge\omega^{2}$ (4.23)
-$\quad+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(d_{\nabla}\alpha\wedge\alpha)\wedge\omega^{2}+\frac{1}{3}\int_{M}\mathop{\mathrm{Tr}}\nolimits(\alpha\wedge\alpha\wedge\alpha)\wedge\omega^{2}.$
-In particular, assuming a transversal deformation, (4.21) holds because the defect term in (4.23) vanishes, as the subintegral quantities in the last two terms belong to $\Omega^{7}_{H}({\mathfrak{g}}_{E})=0$.
-Hence, we can assume that $\alpha$ is a vertical $1$-form, i.e., $\alpha=\eta\otimes s\in\eta\wedge\Omega^{0}_{H}({\mathfrak{g}}_{E})$ for some transversal section $s$, therefore
-$\kappa(\nabla)$ $=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge(\eta\otimes s))\wedge\omega^{2}+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(d_{\nabla}(\eta\otimes s)\wedge(\eta\otimes s))\wedge\omega^{2}$
-$=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\eta\wedge\omega^{2}\otimes s)+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(\omega\otimes s-\eta\otimes\nabla s)\wedge(\eta\otimes s))\wedge\omega^{2}$
-$=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\ast(\omega\otimes s))+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(\eta\wedge\omega^{3}\otimes[s,s]),$
-hence, (4.22) follows. ∎
+Now, using (5.2), we get
+[…]
+$=3g_{im}\,(0,e^{m})=3\,(0,e_{i})=3e_{i}(0,1)$
+$=3e_{i}\cdot\zeta$
+Here, we used the identity $\varphi_{ijk}\varphi_{mnk}=g_{im}g_{jn}-g_{in}g_{jm}+\psi_{ijmn}$ [karigiannis2009flows]*Lemma A.8. Assume $\gamma\in\Lambda^{3}_{27}$, we compute
+$(X\;\lrcorner\;\gamma)\cdot\zeta$ $=X_{i}\,(e_{i}\;\lrcorner\;\gamma)\cdot\zeta=\frac{1}{2}\,X_{i}\,\gamma_{ijk}\,e^{j}\cdot e^{k}\cdot\zeta$
+$=\frac{1}{2}\,X_{i}\bigl(h_{i}{}^{p}\,\varphi_{pjk}+h_{j}{}^{p}\,\varphi_{ipk}+h_{k}{}^{p}\,\varphi_{ijp}\bigr)e^{j}\cdot e^{k}\cdot\zeta.$
+$=\frac{1}{2}\,X_{i}\bigl(h_{i}{}^{p}\,\varphi_{pjk}+2h_{j}{}^{p}\,\varphi_{ipk}\bigr)e^{j}\cdot e^{k}\cdot(1,0)$
+$=\frac{1}{2}\,X_{i}\bigl(h_{i}{}^{p}\,\varphi_{pjk}+2h_{j}{}^{p}\,\varphi_{ipk}\bigr)(0,\,e^{j}\times e^{k})$
+$=\frac{1}{2}\,X_{i}\bigl(h_{i}{}^{p}\,\varphi_{pjk}+2h_{j}{}^{p}\,\varphi_{ipk}\bigr)(0,\,\varphi^{jk}{}_{m}e^{m})$
+$=\frac{1}{2}\,X_{i}\,h_{i}{}^{p}\,(6\,g_{pm})\,(0,e^{m})+X_{i}\,h_{j}{}^{p}\bigl(g_{im}g^{j}{}_{p}-g_{i}{}^{j}g_{pm}+\psi_{ipm}{}^{j}\bigr)(0,e^{m})$
+$=3\,h(X)\cdot\zeta-h(X)\cdot\zeta.$
+$=2\,h(X)\cdot\zeta$
+The next result characterises the infinitesimal deformations of a $\mathrm{G}_{2}$-instanton in terms of the operator (5.7).
+Proposition 5.5.
+Let $\varepsilon\in\Omega^{1}(M,\mathrm{Ad}(P))$ and $D^{t,A}$ the Dirac operator (5.7) with $A$ a $\mathrm{G}_{2}$-instanton.
+Then $\varepsilon$ defines an infinitesimal deformation of AA (cf.
+Definition 5.2), if and only if
+$D^{t,A}(\varepsilon\cdot\zeta)=\frac{(-5+t)}{8}\,\tau_{0}\,\varepsilon\cdot\zeta+(3t-1)\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$
+Proof.
+Let $\{e_{i}\}_{i=1}^{7}$ be a local orthonormal frame on $M$. By definition of the twisted Dirac operator associated with the product connection $\nabla^{0,A}$, a direct computation gives
+$D^{0,A}(\varepsilon\cdot\zeta)$ $=\sum e_{i}\cdot\nabla^{0,A}_{e_{i}}(\varepsilon\cdot\zeta)=(d_{A}\varepsilon+d_{A}^{\ast}\varepsilon)\cdot\zeta+\sum e_{i}\cdot\varepsilon\cdot\nabla^{0,A}_{e_{i}}(\zeta)$
+$=(d_{A}\varepsilon+d_{A}^{*}\varepsilon)\cdot\zeta-\sum_{i}\left(\frac{\tau_{0}}{24}e_{i}\cdot\varepsilon\cdot(e_{i}\lrcorner\varphi)\cdot\zeta-\frac{1}{4}e_{i}\cdot\varepsilon\cdot(e_{i}\lrcorner\tau_{3})\cdot\zeta\right).$
+Here, we used that $X\cdot Y+Y\cdot X=-2g(X,Y)1$ and $\nabla^{1}\zeta=0$ for the connection (5.5). Now, using Lemma 5.4 we get
+$D^{0,A}(\varepsilon\cdot\zeta)$ $=(d_{A}\varepsilon+d_{A}^{*}\varepsilon)\cdot\zeta-\frac{5\tau_{0}}{8}\,\varepsilon\cdot\zeta-(\varepsilon\lrcorner\tau_{27})\cdot\zeta.$
+On the other hand, for $H_{\varphi}=\frac{\tau_{0}}{6}\,\varphi-\tau_{3}$ we have
+[…]
+$=\frac{1}{2}\sum_{i,j,k}H_{ijk}\,\varepsilon_{\ell}\cdot e^{i}\cdot e^{\ell}\cdot e^{j}\cdot e^{k}\cdot\zeta+\sum H_{i\ell k}\varepsilon_{\ell}e^{i}\cdot e^{k}\cdot\zeta-2\,(\varepsilon\;\lrcorner\;H)\cdot\zeta$
+$=-\frac{1}{2}\sum H_{ijk}\varepsilon_{\ell}\cdot e^{\ell}\cdot e^{i}\cdot e^{j}\cdot e^{k}\cdot\zeta-\sum H_{\ell jk}\,\varepsilon_{\ell}e^{j}\cdot e^{k}\cdot\zeta-4\,(\varepsilon\;\lrcorner\;H)\cdot\zeta.$
+$=-3\,\varepsilon\cdot H\cdot\zeta-6\,(\varepsilon\;\lrcorner\;H)\cdot\zeta=-\frac{1}{2}\,\tau_{0}\,\varepsilon\cdot\varphi\cdot\zeta-6\,(\varepsilon\;\lrcorner\;H)\cdot\zeta.$
+$=\frac{7}{2}\,\tau_{0}\,\varepsilon\cdot\zeta-\tau_{0}\,(\varepsilon\;\lrcorner\;\varphi)\cdot\zeta+6\,(\varepsilon\;\lrcorner\;\tau_{3})\cdot\zeta.$
+$=\frac{1}{2}\,\tau_{0}\,\varepsilon\cdot\zeta+12\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$
+Consequently, from (5.7) we obtain
+$D^{t,A}(\varepsilon\cdot\zeta)=(d_{A}\varepsilon+d_{A}^{*}\varepsilon)\cdot\zeta-\frac{5}{8}\,\tau_{0}\,\varepsilon\cdot\zeta-(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta+\frac{t}{8}\,\tau_{0}\,\varepsilon\cdot\zeta+3t\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$
+Assume that $\varepsilon$ satisfies the linearised instanton equation together with the Coulomb gauge condition, i.e., $d_{A}\varepsilon=0$ and $d_{A}^{*}\varepsilon=0$, then
+$D^{t,A}(\varepsilon\cdot\zeta)=\frac{-5+t}{8}\,\tau_{0}\,\varepsilon\cdot\zeta+(3t-1)\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$ (5.8)
+Since $\zeta$ is parallel with respect to $\nabla^{1}$, then $D^{1,A}(\varepsilon\cdot\zeta)=(D^{1,A}\varepsilon)\cdot\zeta,$ in consequence,
+$(D^{1,A}\varepsilon)\cdot\zeta=-\frac{\tau_{0}}{2}\,\varepsilon\cdot\zeta+2\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$ (5.9)
+Hence, the infinitesimal deformation space of a $\mathrm{G}_{2}$-instanton $A$ on a $7$-manifold with coclosed $\mathrm{G}_{2}$-structure is isomorphic to the kernel of the operator
+$D^{1,A}+\frac{1}{2}\,\tau_{0}\,\mathrm{Id}_{\Omega^{1}(M,\mathrm{Ad}(P))}-2\,\tau_{27}^{*}\otimes\mathrm{Id}_{\mathrm{Ad}(P)}\in\mathrm{End}\bigl(\Omega^{1}(M,\mathrm{Ad}(P))\bigr).$ (5.10)
+We have $D^{t,A}=D^{1,A}+(t-1)\,\frac{1}{4}\,g^{-1}\not${H}. Consequently,
+$D^{t,A}(\varepsilon\cdot\zeta)=D^{1,A}(\varepsilon\cdot\zeta)+(t-1)\,\frac{\tau_{0}}{8}\,\varepsilon\cdot\zeta+(t-1)\,3\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$
+Setting $t=\frac{1}{3}$, we obtain $D^{1/3,A}(\varepsilon\cdot\zeta)=D^{1,A}(\varepsilon\cdot\zeta)-\frac{\tau_{0}}{12}\,\varepsilon\cdot\zeta-2\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta$ and consequently,
+$D^{1/3,A}+\frac{7}{12}\,\tau_{0}\,\mathrm{Id}=D^{1,A}+\frac{1}{2}\,\tau_{0}\,\mathrm{Id}-2\,\tau_{27}^{*}\otimes\mathrm{Id}_{\mathrm{Ad}(P)}$ (5.11)
+5.2 Weitzenböck inequality and curvature obstruction
+It is known the Weitzenböck–Lichnerowicz formula with torsion cf. [Agricola2004Holonomy]
+Proposition 5.6.
+[…]
+$(D^{t/3,A})^{2}\mu=(\nabla^{t,A})^{\ast}\nabla^{t,A}\mu+\frac{1}{4}\,\mathrm{Scal}\,\mu+\frac{t}{6}\,d\varphi\cdot\mu-\frac{t^{2}}{18}\,\|\varphi\|^{2}\mu+F\cdot\mu.$ (5.12)
+As an application of Propositions 5.5 and 5.6, we prove the following rigidity result.
 Corollary 5.7.
-Let $E$ be a Sasakian ${\rm G}$-bundle over a compact closed 7-manifold. Any SDCI is a local minimizer of the YM-functional (4.14).
+Let $(V^{5,2},\varphi)$ be the Stiefel manifold endowed with the coclosed $\mathrm{G}_{2}$–structure given in Corollary 2.8 and let AA denote the $\mathrm{G}_{2}$–instanton (4.9) on the ${\rm SO}(3)$-bundle $P_{1}=V^{5,2}\times_{\phi_{1}}{\rm SO}(3)$.
+Then, for $y_{2}>0$ as in (4.12) there exists a non–empty interval $y_{2}\in I$ such that, for all values $y_{2}\in I$, the $\mathrm{G}_{2}$–instanton $A$ is rigid for these values of the metric parameter $y_{2}$.
 Proof.
-From (4.19), note that the YM-functional assumes the value of the charge $\kappa(\nabla)$ (4.17). From (4.22)
-$\kappa(\nabla)$ $=\kappa(\nabla^{\prime})+({\rm F}_{\nabla},\omega\otimes s)_{L^{2}}$
-$=\kappa(\nabla^{\prime}),$
-the last line follows since $\Omega^{2}_{8}({\mathfrak{g}}_{E})\perp\omega\otimes\Omega^{0}({\mathfrak{g}}_{E})$.
-Hence, if $\nabla$ is a SDCI the charge is locally constant.
-The same is true if $\nabla$ is a ASDCI, since $\Omega^{2}_{6}({\mathfrak{g}}_{E})\perp\omega\otimes\Omega^{0}({\mathfrak{g}}_{E})$.
-∎
+[…]
+$\bigl|\nabla^{1/3,A}(\varepsilon\cdot\zeta)\bigr|^{2}$ $=\left\langle(D^{\frac{1}{3},A})^{2}(\varepsilon\cdot\zeta),\,\varepsilon\cdot\zeta\right\rangle-\frac{1}{4}\,\mathrm{Scal}\,|\varepsilon\cdot\zeta|^{2}-\frac{1}{18}\,\langle d\varphi\cdot(\varepsilon\cdot\zeta),\,\varepsilon\cdot\zeta\rangle$ (5.13)
+$\quad+\frac{1}{162}\,\|\varphi\|^{2}|\varepsilon\cdot\zeta|^{2}-\langle F_{A}(\varepsilon\cdot\zeta),\,\varepsilon\cdot\zeta\rangle.$
+From Proposition 5.5, if $\alpha\in\Omega^{1}(M,\mathrm{Ad}(P))$ is an infinitesimal deformation of the $\mathrm{G}_{2}$–instanton $A$, then $\varepsilon$ satisfies $D^{1/3,A}(\varepsilon\cdot\zeta)=-\frac{7}{12}\,\tau_{0}\,\varepsilon\cdot\zeta,$ hence
+$\left\langle(D^{\frac{1}{3},A})^{2}(\varepsilon\cdot\zeta),\,\varepsilon\cdot\zeta\right\rangle=\frac{49}{144}\,\tau_{0}^{2}\,|\varepsilon\cdot\zeta|^{2}.$ (5.14)
+From (2.23), we compute
+$d\varphi$ $=2\left(X^{2356}+X^{2457}+X^{3467}\right)-\frac{3P_{1}}{2y_{2}}\left(X^{1237}-X^{1246}+X^{1345}-X^{1567}\right)$
+$+\frac{3P_{3}}{2y_{2}}\left(X^{1234}-X^{1267}+X^{1357}-X^{1456}\right).$
+where $P_{1}=a(a^{2}+b^{2})^{-1/2}$ and $P_{3}=b(a^{2}+b^{2})^{-1/2}$.
+Notice that $P_{1}^{2}+P_{3}^{2}=1$ and $P_{1}+P_{2}\leq 1$.
+On the other hand, from the proof of Theorem 2.1, we have $\tau_{0}=\frac{6}{7}\left(2+\frac{1}{y_{2}}\right)$ and therefore,
+$\ast\tau_{3}$ $=d\varphi-\tau_{0}\psi$
+$=$ $\left(\frac{26}{7}+\frac{6}{7y_{2}}\right)\left(X^{2356}+X^{2456}+X^{3467}\right)-\left(\frac{6}{7}+\frac{33}{14}\frac{P_{1}}{y_{2}}\right)\left(X^{1237}-X^{1246}+X^{1345}-X^{1567}\right)$
+$+\left(\frac{6}{7}+\frac{33}{14}\frac{P_{3}}{y_{2}}\right)\left(X^{1234}-X^{1267}+X^{1357}-X^{1456}\right).$
+Hence $\lvert\tau_{3}\rvert^{2}=3\left(\frac{26}{7}+\frac{6}{7y_{2}}\right)^{2}+4\left(\frac{6}{7}+\frac{33}{14}\frac{P_{1}}{y_{2}}\right)^{2}+4\left(\frac{6}{7}+\frac{33}{14}\frac{P_{3}}{y_{2}}\right)^{2}$ and using $P_{1}+P_{3}\leq P_{1}^{2}+P_{3}^{2}=1,$ we get
+$\lvert\tau_{3}\rvert^{2}\leq\frac{4}{49}\left(241+276\,\frac{1}{y_{2}}+\frac{1125}{4}\,\frac{1}{y_{2}^{2}}\right).$ (5.15)
+From (5.13), using (5.14)
+$(*)$ $\leq\left(\frac{49\lvert\tau_{0}\rvert^{2}}{144}-\frac{\mathrm{Scal}}{4}-\frac{\lvert\tau_{0}\rvert}{18}+\frac{7}{162}+\frac{1}{18}\lvert\tau_{3}\rvert+\lvert F_{A}\rvert\right)\lvert\varepsilon\cdot\zeta\rvert^{2}.$
+Thus, from (4.12) and (5.15), we obtain
+$0$ $\leq\left(\frac{49\lvert\tau_{0}\rvert^{2}}{144}-\frac{\mathrm{Scal}}{4}-\frac{\lvert\tau_{0}\rvert}{8}+\frac{7}{162}+\frac{\lvert\tau_{3}\rvert}{18}+\lvert F_{A}\rvert\right)\lvert\varepsilon\rvert^{2}$ (5.16)
+$\leq\Bigg[\frac{1}{4}\left(2+\frac{1}{y_{2}}\right)^{2}+3-\frac{6}{y_{2}}-\frac{3}{28}\left(2+\frac{1}{y_{2}}\right)+\frac{7}{162}+\frac{1}{63}\sqrt{241+\frac{276}{y_{2}}+\frac{1125}{4y_{2}^{2}}}+\frac{\sqrt{3}}{y_{2}}\Bigg]\lvert\varepsilon\rvert^{2}.$
+Setting $x=\frac{1}{y_{2}},$ the right–hand side of (5.16) can be written as $f(x)\,|\varepsilon|^{2},$ where for $x\geq 0$
+$f(x)$= $\frac{1}{4}(2+x)^{2}+3-6x-\frac{3}{28}(2+x)+\frac{7}{162}+\frac{1}{63}\sqrt{241+276x+\frac{1125}{4}x^{2}}+\sqrt{3}\,x$
+The function $f$ is continuous on $0,\infty)$, moreover, a direct evaluation shows that $f(x)$ becomes negative on a bounded interval.
+In particular, $f(x)<0\quad\text{for all }x\in(\frac{5}{3},10).$ [^5] This interval is not maximal Equivalently, $f(y_{2})<0$ for all $y_{2}\in\left(\tfrac{1}{10},\tfrac{3}{5}\right)$.
+For such values of the metric parameter $y_{2}$, the right–hand side of the Weitzenböck inequality becomes strictly negative (see Figure 1), while the left–hand side is non–negative.
+This yields a contradiction unless $\varepsilon\equiv 0.$ Therefore, the instanton is rigid for this range of metric parameters.
+xxf(x)f(x)
+Figure 1: Plot of the right–hand side of the Weitzenböck inequality as a function of the metric parameter $y_{2}$.
+This shows that the expression becomes strictly negative for a non-empty interval of positive values of $y_{2}$.